### v1

# A demonstration of the  class file for SAGE Publications

Abstract
Clothes manipulation, such as folding or hanging, is a critical capability for home service robots. Despite recent advances, most existing methods remain limited to specific tasks and clothes types, due to the complex, high-dimensional geometry of clothes. This paper presents CLothes mAnipulation with Semantic keyPoints (CLASP), which aims at general-purpose clothes manipulation over different clothes types, T-shirts, shorts, skirts, long dresses, … , as well as different tasks, folding, flattening, hanging, … . The core idea of CLASP is semantic keypoints—e.g., “left sleeve”, “right shoulder”, etc.—a sparse spatial-semantic representation that is salient for both perception and action. Semantic keypoints of clothes can be reliably extracted from RGB-D images and provide an effective intermediate representation of clothes manipulation policies. CLASP uses semantic keypoints to bridge high-level task planning and low-level action execution. At the high level, it exploits vision language models (VLMs) to predict task plans over the semantic keypoints. At the low level, it executes the plans with the help of a simple pre-built manipulation skill library. Extensive simulation experiments show that CLASP outperforms state-of-the-art baseline methods on multiple tasks across diverse clothes types, demonstrating strong performance and generalization. Further experiments with a Franka dual-arm system on four distinct tasks—folding, flattening, hanging, and placing—confirm CLASP’s performance on a real robot. The CLASP source code is available at https://github.com/dengyh16code/CLASP.
keywords:
Deformable Object Manipulation, State Representation, Foundation Models
1 Introduction
People have long anticipated that an intelligent home service robot would take care of their laundry chores, like folding a T-shirt for storage or hanging a skirt on a hanger. Toward this goal, we need a general-purpose robot capable of manipulating a wide variety of clothes in many different ways. Despite recent significant advances in clothes manipulation skills like folding and flattening (Moletta et al., 2023; Canberk et al., 2023), these methods are tailored to specific clothes types and tasks.
How can we build a robot for general-purpose clothes manipulation? Clothes are deformable objects with a high-dimensional state space, and different types of clothes (e.g., T-shirts and skirts) exhibit distinct geometric structures. This complexity makes it essential to develop an effective and general state representation for general-purpose manipulation. Although physics-based representations, such as particles and meshes (Bender et al., 2015; Chen et al., 2022), can represent a wide range of clothes, they are complex and challenging to estimate accurately, especially under deformation and occlusion. Keypoint-based representation provides a more succinct representation, which uses dense keypoints to capture the topological structure of clothes (Gustavsson et al., 2022; Lips et al., 2024). However, these dense keypoints are also challenging to track and estimate under deformation and occlusion. Additionally, these keypoints are tailored to specific clothes items and do not generalize well across different instances. These limitations arise because previous keypoint representations focus only on geometric information and ignore semantic information, which is crucial given that clothes are manually designed and inherently shaped for humans.
In this paper, we present semantic keypoints as a general spatial-semantic representation of clothes. Compared with previous keypoint representations, semantic keypoints carry explicit semantic meaning and can be intuitively described using natural language. As a result, semantic keypoints offer a more sparse and succinct representation and focus on distinct structural features of clothes like sleeves and shoulders. These structural features are salient for both perception and action. In the context of perception, semantic keypoints are relatively easy to extract from observations and remain consistent across diverse clothes instances of the same type. This consistency is the foundation of generalization. In the context of action, semantic keypoints not only capture the geometric structure of clothes but also identify where clothes are frequently manipulated. As a result, semantic keypoints provide an effective abstraction of the action space for task planning, bridging the gap between perception and manipulation.
Figure 1: CLothes mAnipulation with Semantic keyPoints (CLASP). The semantic keypoint representation enables CLASP to generalize over many different types of clothes and tasks. (a) Semantic keypoints for various types of clothes. (b) Four distinct clothes manipulation tasks in our experiments.
For general-purpose clothes manipulation, we integrate semantic keypoints with foundation models and develop a CLothes mAnipulation method with Semantic keyPoints (CLASP). As Figure 1 shows, CLASP can be deployed on a wide range of clothes types and manipulation tasks. CLASP first leverages foundation models for open-category semantic keypoint extraction. Semantic keypoints are the inherent features of clothes, which can be effectively captured in the commonsense knowledge of foundation models. Specifically, we propose a two-stage pipeline for semantic keypoint extraction. The first stage leverages a vision language model (VLM) for semantic understanding of clothes and autonomously discovers semantic keypoints on a prototype image for each clothes category. The second stage aims to match semantic keypoints on the prototype image to novel clothes, where vision foundation models ensure spatial precision under deformation and occlusion.
To manipulate clothes using semantic keypoints, we first build a skill library. We leverage the commonsense knowledge of a large language model (LLM) to identify general basic skills for clothes manipulation, then implement them as policies parameterized by contact point positions. Given the skill library, CLASP leverages a VLM-powered task planner to decompose the free-form natural language instruction into sequential subtasks conditioned on the observation image and semantic keypoints. Each subtask is described by a basic skill selected from the skill library and contact points like grasp(“left sleeve”). In task planning, semantic keypoints define where clothes can be manipulated and serve as contact point candidates. After obtaining the task plan, CLASP verifies whether the plan is executable. If the task plan fails verification, the failed subtask prompts the VLM to replan; otherwise, CLASP executes the subtasks sequentially. For each subtask, CLASP invokes a policy from the skill library to generate waypoints based on contact point positions and applies motion planning to produce an executable trajectory. After executing each subtask, CLASP updates the observation and determines whether to replan. This closed-loop pipeline will be repeated until the task is completed.
To evaluate CLASP, we extend the SoftGym (Lin et al., 2021) benchmark and conduct simulation experiments across four clothes categories and four representative tasks in various scenarios. CLASP outperforms baseline methods like flingbot (Ha and Song, 2022) and fabricflownet (Weng et al., 2022) in clothes manipulation. Real-world experiments on 15 various clothes show that CLASP can be easily deployed on a real robot, performing well across a wide variety of clothes and tasks. These results demonstrate that semantic keypoints are an effective representation of clothes manipulation, enabling general-purpose clothes manipulation with foundation models.
This paper is an extension of our previous work presented at IEEE International Conference on Robotics and Automation (ICRA), titled “General-purpose Clothes Manipulation with Semantic Keypoints” (Deng and Hsu, 2025). We expand upon the conference version in three primary aspects: (i) Open-category semantic keypoint extraction: we improve the semantic keypoint extraction method by leveraging foundation models, eliminating the need for training on labeled data. This enhancement significantly broadens the applicability and generalization of our method to open-category clothes. (ii) Closed-loop task planning: we improve the previous open-loop system with plan verification, dynamic replanning from new observations, and task completion checks—greatly improving robustness to unexpected state changes, especially with initially crumpled clothes. (iii) Expanded Real-world Evaluations: we extend the scope of real-world experiments to a diverse range, including T-shirts, sweaters, shorts, pants, skirts, long dresses, and more. Additionally, we conducted comprehensive quantitative real-world experiments and analysis to validate the effectiveness and generalization of our method.
2 Related Work
2.1 State Representation for Deformable Object
Effective state representation is crucial for object manipulation. However, representing the state of deformable objects, such as clothes, remains a significant challenge due to their high-dimensional state space (Yin et al., 2021; Longhini et al., 2024a). Some task-specific state representations have been developed, such as using an approximate polygonal representation for clothes folding (Miller et al., 2012) and topological coordinates for assistive dressing (Wang et al., 2023). However, designing a general representation across different objects and tasks remains a significant challenge.
Physics-based models such as particle (Bender et al., 2015; Chen et al., 2022; Zhang et al., 2024b) and mesh (Guler et al., 2015; Macklin et al., 2016) provide a solution to model deformable dynamics and represent a wide range of objects. However, these dense models make accurate state estimation of deformable objects challenging, especially under deformation and occlusion (Lin et al., 2022).
Recently, learning to encode raw observations into a compact latent representation has emerged as a more effective solution (Lippi et al., 2020; Yan et al., 2021). However, they often struggle to generalize to new environments due to distribution shifts caused by variations in visual conditions (Hoque et al., 2022).
Keypoint representation provides a succinct way for deformable object representation, enabling efficient planning and learning (Kulkarni et al., 2019; Ma et al., 2022; Wu et al., 2024) while remaining robust to visual condition variations. Previous keypoints representation focuses on geometric information, using dense keypoints to represent the topology of deformable objects (Shi et al., 2021; Gustavsson et al., 2022; Zhou et al., 2023a). However, keypoint extraction is still challenged by occlusion and deformation, while dense keypoints often lack consistency and generalization. More recently, Huang et al. (2024) utilize vision foundation models to extract sparse keypoints and employs a vision-language model to impose keypoint-centric constraints for clothes manipulation, though it requires domain-specific human intervention.
In this paper, we present semantic keypoints as a general spatial-semantic representation of clothes. Semantic keypoints carry semantic meaning and can be described in natural language, making them sparse and salient for perception and action. Semantic keypoints are easy to extract, generalize well, and indicate where clothes are more likely to be manipulated.
2.2 Deformable Object Manipulation
Existing deformable manipulation approaches can be categorized into model-based and model-free approaches. Model-based approaches use a dynamic model to predict configurations of deformable objects under a certain action and then select the appropriate action (Cusumano-Towner et al., 2011; Miller et al., 2012; Yu et al., 2023). The key challenge lies in the dynamic modeling of deformable objects. To tackle this, two main techniques are used: physics-based models (Li et al., 2018; Chen et al., 2024), which rely on physical law, and data-driven models (Tanaka et al., 2018; Lin et al., 2022), which are learned from interaction data. Typical physics-based models include particle-based models (Longhini et al., 2024b; Zhang et al., 2024b) and mesh-based models (Huang et al., 2022). Recently, Material Point Method (MPM) (Jiang et al., 2016) has emerged as a new physics-based modeling method, which combines particles with grid-based forces for more accurate dynamic modeling (Gu et al., 2023). However, physics-based modes often struggle with heavy parameter tuning (Zheng et al., 2024). Data-driven dynamics models are learned from interaction data, typically formulated as a supervised learning problem (Hu et al., 2018; Tanaka et al., 2018; Yan et al., 2021). Compared to physics-based models, data-driven models support more flexible state space, especially via latent representation (Lippi et al., 2020; Li et al., 2024). However, data-driven models rely on expensive robot interaction data and often fail to generalize to different objects and tasks. Overall, the generalization of model-based approaches highly depends on accurate dynamic modeling and calibration for unseen objects and tasks (Longhini et al., 2024a), which remains an open problem due to the complex, nonlinear dynamics of deformable objects.
Model-free approaches aim to learn robot actions directly without establishing a dynamic model, which can be categorized as imitation learning and reinforcement learning methods. The crucial challenge of imitation learning methods is collecting expert demonstrations, which can be gathered from human (Joshi et al., 2019; Lee et al., 2024) or scripted policies (Jia et al., 2019; Weng et al., 2022; Mo et al., 2022). Specifically, Avigal et al. (2022) employ imitation learning to acquire quasi-static action primitives, such as pick-and-place and dragging, for bimanual cloth folding. Lee et al. (2024) propose a method to learn fabric smoothing and folding policies from human demonstration videos. To avoid collecting extensive expert demonstrations, reinforcement learning methods offer an alternative solution. Matas et al. (2018) develop a sim-to-real reinforcement learning method for towel hanging. Ha and Song (2022) use reinforcement learning to learn a dynamic fling policy for efficient cloth flattening.
Although model-free approaches have made significant progress in a wide range of tasks (Ganapathi et al., 2021; Xu et al., 2022; Bahety et al., 2023), the generalization remains a significant limitation. Most of the previous methods aim to learn task-specific policies (Salhotra et al., 2022; Weng et al., 2022; Gao et al., 2023). Although some goal-conditioned methods use goal images for multi-task learning (Seita et al., 2021; Chi et al., 2022), they often struggle to generalize to unseen goals. To enhance generalization, collecting more data (Xue et al., 2023) and domain randomization (Hietala et al., 2022; Wang et al., 2023) have been explored. However, model-free approaches still struggle to generalize to unseen tasks or object categories. In this paper, we aim to tackle general-purpose cloth manipulation across diverse clothes categories and task types. To this end, we propose CLASP, a framework that integrates semantic keypoint representations with foundation models. CLASP enables generalization to a wide range of clothes and manipulation tasks.
Figure 2: CLASP Overview. Given an RGB-D observation, CLASP extracts semantic keypoints as the state. These keypoints, along with the RGB image and task instruction, are fed to a VLM to generate a task plan. Once the plan is verified, the subtasks are executed sequentially. After each subtask, CLASP updates the observation and determines whether to replan. This process repeats until the overall task is completed.
2.3 Language-Conditioned Manipulation
Language-conditioned manipulation is an emerging field aimed at enabling seamless interaction between humans and robots by teaching robots to understand and execute instructions specified in natural language (Shridhar et al., 2022; Mees et al., 2022; Zhou et al., 2023b). The recent emergence of large foundation models, such as large language models (LLMs) and vision-language models (VLMs), has significantly advanced language-conditioned manipulation tasks. These models provide commonsense knowledge for reasoning (Singh et al., 2023; Du et al., 2024) and perception (Huang et al., 2023; Gao et al., 2024), significantly enhancing generalization capabilities.
Beyond specifying tasks, natural language also serves as a powerful abstraction of affordance (Mees et al., 2023), capturing the functions of an object, i.e., how it can be manipulated. For task-oriented grasping (Liu et al., 2020), GraspGPT (Tang et al., 2023) utilizes object class descriptions generated by a large language model (LLM) to identify optimal grasping regions. Similarly,  Ren et al. (2023) provide natural language descriptions of tools to accelerate learning language-conditioned tool manipulation policies. RoboPoint fine-tunes a VLM for spatial affordance prediction conditioned on natural language (Yuan et al., 2024).
Most previous language-conditioned object manipulation methods focus on rigid objects, with few addressing deformable object manipulation (Deng et al., 2024; Barbany et al., 2025). For rigid objects, existing methods are available to execute actions such as picking, placing, moving, and grasping (Liang et al., 2023). However, a significant gap remains between task planning and action execution when it comes to deformable object manipulation. Motivated by prior work that uses natural language as an abstraction of affordance, we propose semantic keypoints as a spatial-semantic representation of clothes. These keypoints act as an interface between VLM-powered task planning and low-level action execution, enabling general-purpose clothes manipulation.
3 Overview
In this paper, we introduce CLASP (CLothes mAnipulation method with Semantic keyPoints), a general-purpose clothes manipulation method. The core idea of CLASP is to use semantic keypoints as a general state representation for clothes. Each semantic keypoint includes: (i) a language description (e.g., “left sleeve”), providing semantic context for high-level task planning, and (ii) a keypoint position, offering geometric information to guide low-level action generation.
To leverage semantic keypoints in clothes manipulation, we first build a basic skill library. Instead of manually designing skills, we prompt an LLM to discover basic clothes manipulation skills. The LLM’s commonsense knowledge ensures these skills are general and sufficient. We then implement the discovered skills as policies parameterized by contact point positions, to form the skill library.
Fig. 2 illustrates the overall framework of CLASP. Given RGB-D observations of clothes and a language instruction, CLASP generates sequential manipulation actions to complete the task specified by the instruction. CLASP first extracts semantic keypoints through foundation models. After extracting semantic keypoints, the RGB image marked with semantic keypoints and the language instruction are fed into a VLM for task planning. The VLM determines whether the task is complete and infers a task plan, where each subtask consists of a basic skill and contact points. After obtaining the task plan, CLASP verifies whether the plan is executable. The verified task plan will be executed sequentially by invoking policies from the skill library. After each subtask, CLASP updates the observation and decides whether to replan. The closed-loop pipeline repeats until the task is completed.
In the following subsections, we will introduce how we extract effective semantic keypoints of clothes in Sec. 4. Then, to enable general-purpose clothes manipulation with semantic keypoints, we present the construction of the skill library in Sec. 5.1, followed by the details of task planning and execution in Sec 5.2.
4 Semantic Keypoints
Although previous work has explored using keypoints for clothes manipulation, most methods rely on manually annotated keypoints (Mo et al., 2022; Lips et al., 2024). Since clothes are highly deformable, collecting sufficient data that captures clothes’ various configurations is time-consuming and labor-intensive. To mitigate data scarcity, augmenting 3D synthetic data in the simulation environment has been explored (Deng and Hsu, 2025). However, this approach is difficult to scale to open-category clothes due to the limited availability of 3D clothes models. Recently, researchers have explored zero-shot keypoint extraction by leveraging powerful and generalizable feature maps from vision foundation models (Huang et al., 2024; Di Palo and Johns, 2024; Fang et al., 2024b). However, these methods can not ensure that the resulting keypoints have explicit semantic meaning and can be described using natural language. In addition, using the vision foundation model for consistent keypoint extraction on clothes remains challenging, especially under self-occlusion and deformation.
Given that semantic keypoints encode both semantic and spatial information, two key technical challenges must be addressed in keypoint extraction: understanding the semantics of open-category clothes and maintaining the spatial precision of keypoints under deformation and occlusion. To address them, we propose a two-stage pipeline for keypoint extraction. The first stage (Fig. 3) focuses on semantic understanding, aiming to autonomously discover semantic keypoints on prototype images and describe them using natural language. Specifically, we leverage geometric prior and the vision language model’s commonsense knowledge to develop an autonomous semantic keypoint discovery pipeline. For each clothes category, we provide a prototype image free from occlusion and deformation, which makes semantic understanding feasible and also provides full observation as a reference in the second stage. The second stage (Fig. 5) focuses on spatial precision, aiming to match semantic keypoints on the prototype image to novel clothes, which may involve occlusions and deformations. Specifically, we utilize a coarse-to-fine matching strategy to propose semantic keypoints on novel clothes precisely.
Figure 3: Semantic Keypoint Detection. We design an autonomous pipeline to discover semantic keypoints on a prototype image without occlusion and deformation for each clothes type.
Figure 4: Prototype Image Retrieval. Given an RGB observation image of novel clothes, we retrieve the most relevant prototype image by computing Best Buddies Similarity (BBS) using DINOv2 features.
Figure 5: Semantic Keypoint Matching. Given the prototype image and discovered semantic keypoints, we matched the discovered semantic keypoints to novel clothes with VLM and vision foundation models.
4.1 Semantic Keypoint Discovery
For each clothes type, we provide a prototype image free from occlusion and deformation, allowing the semantic keypoint discovery stage to focus on understanding semantics and defining semantic keypoints for each clothes type. To discover semantic keypoints on the prototype image, we design an autonomous pipeline as shown in Figure 3, which leverages the geometric prior of clothes and the semantic understanding capability of the VLM for keypoint discovery.
To discover semantic keypoints, we first extract the contour and skeleton of clothes, which serve as an effective abstraction, capturing the geometric structure of clothes and encoding the human body. Specifically, we employ OWL-V2 (Minderer et al., 2024) to generate the bounding box of clothes on the prototype image based on the text prompt. Later, the bounding box will serve as the box prompts for the Segment-Anything Model (SAM) (Kirillov et al., 2023) to generate the corresponding masked image $I_{m}$ of the clothes. We then extract the contour of clothes from the masked image $I_{m}$ and perform skeletonization by calculating equidistance to the boundaries (Lee et al., 1994), resulting in the skeleton. To get keypoints from the skeleton and boundary, we then apply farthest point sampling (fps) (Eldar et al., 1997) on the boundary and skeleton to sample candidate keypoints $\boldsymbol{K}_{fps}$. Farthest point sampling is an effective method for obtaining diverse, spatially distributed candidate keypoints that are geometrically salient.
After obtaining candidate keypoints $\boldsymbol{K}_{fps}$, selecting semantic keypoints requires semantic understanding of clothes and clothes manipulation. Therefore, we leverage the commonsense knowledge of the VLM and employ mark-based visual prompting (Fang et al., 2024a) to select semantic keypoints and describe them with natural language. In this paper, we use GPT-4o as the VLM. The prompt is multi-modal and consists of two parts: (i) text prompt, and (ii) marked image. In the text prompt, we provide context information to the VLM, including the definition of semantic keypoints, the expected behavior of the VLM, and the desired response format. In the marked image, we draw visual marks on candidate keypoints $\boldsymbol{K}_{fps}$ and label them with numbers. Through mark-based visual prompting, we can get the semantic keypoints $\boldsymbol{K}_{sem}=\{p_{i},s_{i}\}_{i=1}^{n}$, where $p_{i}$, $s_{i}$ refer to the pixel position and language description of $i$-th semantic keypoint, respectively, and $n$ represents the total number of semantic keypoints.
In this way, we establish a prototype set, where each image represents a clothes type and is annotated with semantic keypoints.
4.2 Semantic Keypoint Matching
Although the autonomous pipeline can discover semantic keypoints on the prototype image in the first stage, it does not perform robustly on clothes with occlusion and deformation, which may lead to failures in boundary and skeleton extraction. Thus, the second stage performs semantic keypoint matching to transfer the semantic keypoints from the prototype image to novel clothes. The second stage ensures precise semantic keypoints extraction under occlusion and deformation.
Given an RGB observation image of novel clothes, we first retrieve the most relevant prototype image by computing the Best Buddies Similarity (BBS) using features extracted from the vision foundation model DINOv2 (Oquab et al., 2023). BBS is a robust method for measuring structural similarity between two images. It identifies pairs of local patches—one from the observation image and one from the prototype image—that are mutual nearest neighbors in the feature space. These pairs, referred to as best buddy pairs, capture strong local correspondences. The BBS score is defined as the proportion of local patches in the observation image that form such pairs. By evaluating the BBS score across all prototype images, we retrieve the one that is most structurally similar to the observation image. To enhance robustness, as shown in Figure 4, we apply a set of rotations to the observation image and select the rotation that yields the highest BBS score.
If the highest BBS score within the prototype image set falls below a predefined threshold, the observation is classified as a new clothes type, requiring a new prototype image. This new prototype image is then annotated using our autonomous semantic keypoint discovery pipeline. Otherwise, the prototype image with the highest BBS score is retrieved.
After retrieving the most relevant prototype image $I_{r}$ annotated with semantic keypoints, we then perform a coarse-to-fine semantic keypoint matching on novel clothes. Given the novel clothes’ RGB observation image $I_{q}$ and a semantic keypoint position $p_{i}$ on the prototype image, we first match $p_{i}$ to a coarse region $\Omega$ on the observation image $I_{q}$. This stage is crucial for capturing global context and avoiding local minima. Specifically, we employ VLM with mark-based visual prompting for coarse region proposals. We first mark the observation image $I_{q}$ with an overlaid grid, where each grid cell is indexed by a unique text label, and mark the prototype image $I_{r}$ with a dot on $p_{i}$. We then prompt the VLM with the marked observation and prototype image to select the grid indices corresponding to the prototype semantic keypoint position $p_{i}$. To get precise semantic keypoint positions, we then perform fine-grained keypoint matching on the region $\Omega$ with vision foundation models DINOv2 (Oquab et al., 2023) and stable diffusion (Rombach et al., 2022). Borrowing the method from (Zhang et al., 2024a), we feed $I_{r}$ and $I_{q}$ into vision foundation models, respectively, and get feature maps $\mathcal{F}_{r}$ and $\mathcal{F}_{q}$. In this way, we can get the feature $\mathcal{F}_{q}(p)$ of each point $p$ on the region $\Omega$ and the prototype semantic keypoint’s feature $\mathcal{F}_{r}(p_{i})$. We then calculate the similarity score between them to achieve fine-grained keypoint matching:
Figure 6: Skill Library. Our skill library consists of 5 basic skills: grasp, moveto, release, rotate and pull.
$p_{i}^{\prime}=\arg\max_{p\in\Omega}\langle\mathcal{F}_{q}(p),\mathcal{F}_{r}(p_{i})\rangle$ (1)
[…]
5 Clothes Manipulation With Semantic Keypoints
5.1 Skill Library
For general-purpose clothes manipulation, we build a skill library at first. To ensure that the skill library is general and sufficient for various clothes and manipulation tasks, we first prompt the LLM to propose basic skills of clothes manipulation with the chain of thought prompting (Wei et al., 2022). Specifically, the LLM is prompted to (i) generate examples of clothes manipulation tasks, (ii) break down these tasks into basic actions, and (iii) summarize the actions identified in step (iii) to derive a set of basic skills. In this way, we identify the basic skill set $\mathcal{A}$, including grasp: pick up a part of the clothes, moveto: transport the clothes to a target position, release: drop a part of the clothes, rotate: rotate the clothes, and pull: stretch the clothes for flattening. Figure 6 visualizes the basic skills discovered by the LLM. Since semantic keypoints serve as the effective affordances for clothes manipulation and define where clothes are more likely to be manipulated, we implement these basic skills as policies parameterized by contact point positions. Specifically, we define heuristic rules to create waypoints based on the contact point positions for each basic skill. Once the waypoints are determined, a motion planning algorithm is applied to generate the complete trajectory for the robot’s execution. Specifically, we implement each basic skill as follows:
[…]
release()()( ): The skill release drops clothes parts by opening grippers. After the release action, the observation is updated, and CLASP extracts new semantic keypoints.
[…]
5.2 Task Planning
Figure 7: VLM-powered Task Planner. We implement a closed-loop VLM-powered task planner capable of dynamic replanning based on observations, overall task completion checking, and plan verification.
After establishing the skill library for clothes manipulation, we employ a VLM for task planning based on semantic keypoints, taking advantage of its extensive commonsense knowledge (Zhao et al., 2024) gained from internet-scale data. To generate the task plan, we leverage the multi-modal reasoning capabilities of the VLM and employ visual prompting (Figure 7). Specifically, the prompt includes a meta-prompt and few-shot examples. The meta-prompt defines the task objective, describes the available basic skills, and specifies the input-output format. To improve the robustness of task planning, we leverage in-context learning by providing the VLM with few-shot examples consisting of language instructions, an observation image annotated with extracted semantic keypoints, and the corresponding task plan. These examples not only illustrate the expected input-output format but also demonstrate how to utilize the basic skills from the predefined skill library.
Algorithm 1 Task Planning and Verification
0: Language instruction $\mathcal{L}$, observation image $I_{q}$, semantic keypoints $\boldsymbol{K}_{sem}^{\prime}$, skill set $\mathcal{A}$, max attempts $N_{\max}$
0: Verified task plan or failure reason
1:Step 1: Task Completion check
[…]
3:  return  “Task Completed”
4:end if
[…]
6:repeat
7:  $n_{\text{plan}}\leftarrow n_{\text{plan}}+1$
8:  Step 2: Task Plan generation
[…]
10:  Step 3: Plan Verification
11:  failure_reason $\leftarrow$ None
12:  for $\boldsymbol{w}_{j}=(a_{j},c_{j}^{1},c_{j}^{2})$ in $\{\boldsymbol{w}_{j}\}$ do
13:   if $a_{j}\notin\mathcal{A}$ or $c_{j}^{1,2}\notin\boldsymbol{K}_{sem}^{\prime}$ then
14:    failure_reason $\leftarrow$ “Invalid subtask $\boldsymbol{w}_{j}$
15:    break
[…]
17:    failure_reason $\leftarrow$ “Infeasible motion in $\boldsymbol{w}_{j}$
18:    break
19:   end if
20:  end for
21:until failure_reason = None or $n_{\text{plan}}=N_{\max}$
22:if failure_reason = None then
23:  return  $\{\boldsymbol{w}_{j}\}_{j=1}^{H}$
24:else
25:  return  failure_reason
26:end if
The task planning procedure is described in Algorithm LABEL:alg:task_planning. Given a language instruction $\mathcal{L}$, observation image $I_{q}$, and extracted semantic keypoints $\boldsymbol{K}_{sem}^{\prime}$, the VLM decides whether the task is completed and generates task plan $\{\boldsymbol{w}_{j}\}_{j=1}^{H}$, where $H$ denotes the number of subtasks. Each subtask consists of a basic skill and two contact points for the dual-arm manipulation system: $\boldsymbol{w}_{j}=(a_{j},c_{j}^{1},c_{j}^{2})$. The basic skill $a_{j}$ is selected from the predefined set $\mathcal{A}$ introduced in Sec 5.1. The contact points $c_{j}^{1}$ and $c_{j}^{2}$ are selected from extracted semantic keypoints $\boldsymbol{K}_{sem}^{\prime}$. Semantic keypoints define where clothes can be manipulated and serve as contact point candidates in high-level task planning. It should be noted that both $c_{j}^{1}$ and $c_{j}^{2}$ are optional, depending on the number of required contact points. For example, no contact point is required for the skill release. After obtaining the task plan, CLASP proceeds with plan verification by sequentially validating each subtask. For each subtask, CLASP first performs syntax and semantic analysis to ensure that the invoked skill is defined within the skill library and contact points correspond to valid semantic keypoints. It then verifies the subtask at the motion level. Specifically, CLASP invokes a policy from the skill library to create waypoints and perform motion planning to generate the trajectory. The subtask is verified if it passes motion planning. If a subtask fails verification, it is reported as the failure reason to the VLM, prompting the generation of a new task plan.
After obtaining the verified task plan, CLASP executes the subtasks sequentially. Following the completion of each subtask, the RGB-D observation is updated. However, due to occlusion caused by the robot, semantic keypoints are only extracted and updated after the release action. Once the semantic keypoints are updated, the VLM is prompted to replan based on the new observation and updated keypoints. The above process repeats until the task is completed.
6 Experiments
Figure 8: Qualitative Results of Semantic Keypoint Extraction. Our method can achieve open-category semantic keypoint extraction under irregular deformation and occlusion.
Our experiments investigate three questions:
Q1. Can CLASP’s submodules perform effectively under varying conditions?
To answer this, we evaluate the performance of CLASP’s submodules—semantic keypoint extraction and task planning. Results in Sec. 6.1 show that our method can extract effective semantic keypoints on a wide range of clothes even under occlusion and deformation. Results in Sec. 6.2 show that our task planning method can generate reliable and executable task plans.
Q2. How well does CLASP perform on clothes manipulation compared with the existing baseline method?
We address this question through comparative experiments detailed in Sec. 6.3. CLASP outperforms baseline methods in terms of manipulation success rate and generalization capability, highlighting its effectiveness.
Q3. Can CLASP generalize to a wide range of clothes types and manipulation tasks in the real world?
To answer this, we conduct extensive real-world evaluations, as presented in Sec. 6.4. Results show that CLASP generalizes well to various clothes types and manipulation tasks, confirming its practical applicability.
6.1 Semantic Keypoint Extraction Experiments
Table 1: Semantic Keypoint Extraction Results.
Expert AKD (pixel) $\textit{AP}_{15}$ (%) $\textit{AP}_{30}$ (%) $\textit{AP}_{45}$ (%)
1 16.2 71.0 88.0 90.7
2 17.8 69.0 84.1 90.4
3 19.7 68.7 82.2 87.8
We first evaluated the performance of CLASP in extracting effective semantic keypoints. Since no groundtruth semantic keypoints are available, we provide annotated prototype images to three human experts and ask them to label semantic keypoints on novel clothes as ground truth. In this way, we establish a testing set of 120 images from 15 different clothes. We then compare the semantic keypoints extracted by CLASP against those labeled by the experts. In these images, clothes have diverse configurations, poses, occlusions, and deformations.
The evaluation utilizes two standard metrics for keypoint extraction: Average Keypoint Distance (AKD) and Average Precision (AP). AKD calculates the average distance between the extracted keypoints and the groundtruth, while AP measures the proportion of correctly extracted keypoints under different thresholds. Since the resolution of the observed RGB-D images is 720×1280, we used thresholds of 15, 30, and 45 pixels for AP calculation.
The results are shown in TABLE 1. The average keypoint distance between the semantic keypoints proposed by human experts and our method is less than 20 pixels, while the average precision under the 30 pixels threshold is more than 80%.
Considering the resolution and uncertainty in expert labeling, such as identifying the center of clothes, the results can demonstrate that our proposed method performs well in extracting semantic keypoints. For each clothes type, we only need to provide one prototype image without any annotation. Figure 8 shows some qualitative results of semantic keypoint extraction. The experiment results demonstrate that our two-stage semantic keypoint extraction pipeline is effective in extracting semantic keypoints of clothes even under occlusion and deformation. The prior knowledge from the vision language model and vision foundation models ensures the generalization to diverse clothes types. Besides, coarse-to-fine keypoint matching ensures spatial precision and makes semantic keypoint extraction robust.
Table 2: Task planning evaluation results.
Task Check Completion (%) Subtask Accuracy (%)
VLM VLM + KP LLM VLM VLM + KP
Folding 66.7 93.3 33.3 56.7 86.7
Flattening 73.3 96.7 50.0 60.0 93.3
Hanging 93.3 100.0 96.7 76.7 100.0
Placing 83.3 96.7 96.7 70.0 96.7
6.2 Task Planning Experiments
To evaluate the performance of the proposed task planning pipeline for clothes manipulation, we collect images annotated with ground-truth semantic keypoints (labeled by human experts) along with free-form natural language descriptions to evaluate the pipeline on folding, flattening, hanging, and placing tasks. For each task category, we collect 30 images of 15 diverse clothes. We evaluate the pipeline using two metrics: check completion accuracy, which assesses whether the pipeline correctly determines if the overall task has been completed based on the observation; and subtask accuracy, which measures whether the generated subtask sequence is appropriate for the observed state. To further analyze the role of semantic keypoints in task planning, we compare our pipeline (VLM + KP) with two baselines: (i) a VLM baseline that takes raw images without semantic keypoints, and (ii) an LLM baseline that receives only the language descriptors of semantic keypoints along with the language instruction, without any visual input. All pipelines receive the same few-shot demonstrations, but differ in visual input: our method uses keypoint-annotated images, the VLM baseline uses raw images, and the LLM baseline doesn’t rely on visual input.
Table 2 presents the experimental results. Since the LLM baseline lacks visual input, we don’t report its check completion accuracy. It performs the worst on folding and flattening tasks, where adaptation to varied initial states is crucial. For the folding task, some parts of the clothes (such as sleeves) may already be folded. For flattening, the model needs to understand the clothes’ state from visual input and determine whether flattening is required. For example, in some flattening tasks, the robot must repeatedly flatten the clothes to achieve a flat state. The LLM baseline shows slightly better performance in hanging and placing tasks, which involve less uncertainty in the initial state. Moreover, the LLM’s symbolic reasoning remains consistent, robust, and unaffected by distribution shifts in visual input.
Our pipeline achieves reliable accuracy in proposing appropriate task plans and determining task completion. By providing semantic keypoints and a predefined skill library, we enable structured and controllable planning. To enhance robustness, we employ in-context learning with examples that include language instructions, observation images with semantic keypoint annotations, and corresponding task plans. All generated plans are also verified before execution.
Compared to the VLM baseline without semantic keypoints, our pipeline achieves more reliable task planning, especially in long-horizon folding tasks. Without semantic keypoints, the VLM often hallucinates or generates wrong plans. Semantic keypoints provide critical structural information about the clothes’ state, especially under distribution shifts between observation images and the few-shot examples in the prompt. This allows the model to plan based on clothes’ state, rather than being misled by irrelevant visual differences.
Table 3: Simulation Experiment Results. The average success rates (%) on testing tasks. The best performance is in bold.
Method
Folding
(seen object)
Flattening
(seen object)
Hanging
(seen object)
Placing
(seen object)
Towel T-shirt T-shirt Skirt Trousers Towel Towel Skirt
CLIPORT (Shridhar et al., 2022) 77.5 80.0 32.5 36.7 76.7 83.3 93.3 93.3
Goal-conditioned Transporter (Seita et al., 2021) 83.3 76.7 26.7 33.3 100.0 80.0 66.7 83.3
FlingBot (Ha and Song, 2022) N/A N/A 66.7 85.0 N/A N/A N/A N/A
FabricFlowNet (Weng et al., 2022) 93.7 100.0 N/A N/A N/A N/A N/A N/A
CLASP 100.0 95.0 65.0 80.0 96.7 97.5 96.7 96.7
Method
Folding
(unseen object)
Flattening
(unseen object)
Hanging
(unseen object)
Placing
(unseen object)
Trousers Skirt Trousers Towel T-shirt Skirt Trousers T-shirt
CLIPORT (Shridhar et al., 2022) 0.0 0.0 8.3 9.2 76.7 66.7 70.0 73.3
Goal-conditioned Transporter (Seita et al., 2021) 0.0 0.0 10.0 6.7 36.7 40.0 36.7 60.0
FlingBot (Ha and Song, 2022) N/A N/A 29.2 34.2 N/A N/A N/A N/A
FabricFlowNet (Weng et al., 2022) 0.0 2.5 N/A N/A N/A N/A N/A N/A
CLASP 87.5 81.7 60.8 65.0 93.3 76.7 93.3 94.2
6.3 Simulation Experiments
To evaluate CLASP’s performance on clothes manipulation tasks, we conduct experiments in SoftGym, where clothes are modeled as particles with ground truth positions. We sample 3D meshes of clothes from CLOTH3D (Bertiche et al., 2020) dataset, covering four common clothes types in human’s daily life: T-shirts, trousers, skirts, and towel. For each type, over 35 instances of varying sizes and shapes are used. In addition, we extend the SoftGym benchmark to 16 clothes tasks. These tasks can be divided into 4 types:
  • Folding tasks involve folding clothes to achieve a goal configuration. The success of a folding task is determined by the mean particle position error between the achieved and goal configuration.
  • Flattening tasks involve flattening crumpled clothes with random deformations and occlusion. The success of a flattening task is determined by the coverage area of the clothes.
  • Hanging tasks require hanging clothes on a rack. A hanging task is successful when the clothes are fully hung on the rack without any part touching the ground.
  • Placing tasks require placing the clothes in a box. A placing task is successful when the clothes are laid flat inside the box.
Baseline methods include two general multi-task learning frameworks (CLIPORT and Goal-conditioned Transporter) and two task-specific algorithms (FlingBot and FabricFlowNet):
  • CLIPORT (Shridhar et al., 2022) represents a typical end-to-end algorithm for learning language-conditioned manipulation policies, which leverages a pretrained vision language model and a two-stream architecture for semantic understanding and spatial precision.
  • Goal-conditioned Transporter (Seita et al., 2021) represents a typical goal-conditioned transporter network for deformable object manipulation, which infers the optimal action from the current and goal images with a TransporterNets structure.
  • FlingBot (Ha and Song, 2022) is a self-supervised learning framework designed for flattening crumpled clothes by using an effective fling action.
  • FabricFlowNet (Weng et al., 2022) is a learning framework designed for clothes folding, which infers the optimal action based on optical flow.
We also implement an open-loop version of CLASP, where an LLM is prompted to generate a task plan based on the language instruction and semantic keypoint descriptors. In this version, task planning is performed only once at the beginning of the manipulation task and is not conditioned on the observation image.
For each baseline method, we provide 1000 expert demonstrations per task for model training. To evaluate the generalization of baseline methods, only half of the tasks are seen during training through expert demonstrations. For CLASP, we provide few-shot examples of seen tasks in the prompt. Unseen tasks involve new object types. For each task, we conduct 120 trials with different clothes configurations to calculate the success rate. Two task-specific baseline methods are designed for clothes folding and flattening, respectively. We only test them on related tasks, and other tasks are marked as N/A (Not Applicable).
The experiment results are shown in TABLE 3. Overall, CLASP outperforms the two multi-task learning methods on both seen and unseen tasks. Compared with Goal-conditioned Transporter, CLIPORT shows better generalization. Language instructions facilitate capturing similarities between different tasks at the symbolic level. Besides, the pretrained vision-language model enables CLIPORT to capture such similarities. However, CLIPORT’s generalization is limited when adapting to long-horizon tasks, such as transferring the skill from folding a T-shirt to folding trousers, because it learns task-specific manipulation action sequences in an end-to-end manner, and the learned action sequences are not transferable. In contrast, CLASP learns task-agnostic and generalizable language and visual concepts across a wide range of clothes manipulation tasks. The commonsense knowledge from VLM allows CLASP to handle clothes manipulation tasks by decomposing them into predefined basic skills. Furthermore, semantic keypoints are a general spatial-semantic representation of clothes, which are task-agnostic and provide effective cues for task planning and action generation in clothes manipulation.
Compared to the two task-specific algorithms, CLASP shows comparable performance on seen clothes folding and flattening tasks, demonstrating the effectiveness of the proposed method for clothes manipulation. However, on unseen clothes types, two task-specific algorithms tend to fail due to distribution shifts in observations. In contrast, CLASP is a general-purpose clothes manipulation method.
Figure 9: Real World Setup. We establish a dual-arm robot manipulation system for conducting real-world experiments on diverse clothes types.
Table 4: Real Experiment Results. The average success rates on various clothes and tasks.
Task Clothes ID
1 2 3 4 5 6 7 8 9 10 11 12 13 14 15 Mean
Folding 4/5 5/5 4/5 4/5 5/5 5/5 4/5 5/5 3/5 4/5 4/5 5/5 3/5 5/5 4/5 4.3/5
Flattening 2/5 4/5 4/5 3/5 3/5 4/5 2/5 5/5 2/5 3/5 3/5 4/5 3/5 3/5 4/5 3.3/5
Hanging 5/5 5/5 5/5 5/5 N/A 5/5 4/5 5/5 5/5 4/5 4/5 5/5 N/A N/A N/A 4.7/5
Placing 4/5 5/5 5/5 5/5 N/A 5/5 5/5 5/5 5/5 4/5 3/5 5/5 N/A N/A N/A 4.6/5
6.4 Real-world Experiments
To evaluate the performance and generalization of CLASP in real-world scenarios, we establish a dual-arm robot manipulation system. As shown in Figure. 9, the system consists of two Franka Research 3 robot arms and a top-down RealSense L515 RGB-D camera for capturing RGB-D images. The clothes are placed on a table in front of the robot. We use SAM (Kirillov et al., 2023) and OWLv2 (Minderer et al., 2024) to segment the clothes for semantic keypoint extraction. To generate the dual-arm trajectories, we implement a motion planning algorithm using MoveIt! (Chitta et al., 2012) to avoid collisions and synchronize the motion between the arms.
In real-world experiments, we evaluate CLASP on 15 clothes across diverse types, such as shirts, T-shirts, sweaters, skirts, long dresses, and vests, size (ranging from 18 cm × 30 cm for baby pants to 42 cm × 101 cm for adult long dresses), shapes (long sleeves, short sleeves, and sleeveless), and materials (e.g., cotton, polyester, and wool). We directly evaluate CLASP’s performance on manipulation tasks of these clothes. Each clothes item is evaluated across four tasks: folding, flattening, hanging, and placing. The evaluation process includes five trials per task, with success rates calculated as follows. Folding: Success is determined by the Intersection over Union (IoU) between cloth masks achieved by CLASP and human experts. If the IoU is 90% or higher, the task is considered successful. Flattening: if the flattened clothes cover at least 90% of their maximum possible area in a fully flat state, the flattening task is successful. Hanging: the task is considered successful if the clothes are completely hung on the rack without any part touching the ground. Placing: the task is considered successful when the clothes are entirely placed within the box.
The experiment results are shown in Table 4. For certain clothes (#5, #13, #14, and #15), due to limitations in the robot’s working space, we did not perform evaluations on hanging and placing tasks. These are marked as N/A (Not Applicable). Overall, CLASP can be applied to a wide variety of clothes and manipulation tasks while achieving a high success rate. CLASP achieves an 86% success rate in clothes folding, a 66% success rate in clothes flattening, a 94% success rate in clothes hanging, and a 92% success rate in clothes placing. The success rate is comparable to existing task-specific clothes manipulation algorithms, while we test CLASP on broader clothes types and instances. With a general semantic keypoints representation and the prior knowledge from foundation models, CLASP is an effective general-purpose clothes manipulation method. Fig. 10 illustrates some representative examples. For each task, we visualize the extracted semantic keypoints, manipulation action sequences, and achieved object state. CLASP demonstrates the ability to manipulate various clothes in different ways in real-world scenarios. More videos of real-world experiments can be found in our supplementary material.
Figure 10: Qualitative Results of Real-world Experiments. From left to right, the figure displays the task description, extracted semantic keypoints, the real robot’s execution process, and the final achieved state.
7 Discussion
Despite the promising experimental results, CLASP has several limitations: (1) Semantic keypoint extraction under significant occlusion: CLASP leverages existing vision foundation models and the VLM to extract semantic keypoints on open-category clothes without additional training. These models struggle to handle clothes with significant occlusion due to the limited representation of such data in their training datasets. Moreover, inferring the geometric structure of clothes under significant occlusion requires the model to possess a strong prior. This limitation contributes to the relatively lower success rate of CLASP in the clothes flattening task compared to other tasks. State estimation of deformable objects under significant occlusion remains an open challenge in robot perception. One potential improvement could involve designing heuristic policies to reduce the self-occlusion of clothes. This could involve identifying and properly separating different clothes layers or utilizing tools like Breeze to reduce occlusion (Xu et al., 2022). (2) Large clothes manipulation: CLASP struggles to generate executable subtasks for manipulating large clothes due to workspace constraints. This limitation could be addressed by integrating a mobile manipulator system (Yang et al., 2024) or using tools like a hook to pull distant clothes into the workspace (Toussaint et al., 2018). (3) Action execution: During the clothes manipulation process, sliding or deformation may influence the results. In CLASP, semantic keypoint extraction is based on large vision foundation models, making real-time semantic keypoint extraction impossible. One possible improvement could involve using distillation to train a smaller keypoint extraction model, which can track semantic keypoints during clothes manipulation and enable online trajectory optimization.
8 Conclusion
In this paper, we present semantic keypoints as a general spatial-semantic representation of clothes. Building on this representation, we propose CLothes mAnipulation with Semantic keyPoints (CLASP), where semantic keypoints play a crucial role in both high-level task planning and low-level action generation. By integrating the general semantic keypoint representation with foundation models pretrained on internet-scale data, CLASP serves as a general-purpose method for clothes manipulation. Simulation experiments demonstrate that our method outperforms baseline methods in terms of success rate and generalization in clothes manipulation. Real-world experiments further validate CLASP’s generalization in clothes manipulation. CLASP can be directly applied to a diverse range of clothes types and manipulation tasks in real-world scenarios.
Appendix A Discussion of Differences with the Conference Version
This manuscript is an extension of our previous work presented at the IEEE International Conference on Robotics and Automation (ICRA), titled “General-purpose Clothes Manipulation with Semantic Keypoints” (Deng and Hsu, 2025). We have significantly expanded upon the conference version in three primary aspects:
  • 1.
    Introduction of a new open-category semantic keypoint extraction pipeline, enhancing applicability and generalization (Section 4).
  • 2.
    Introduction of a closed-loop task planning pipeline capable of plan verification, dynamic replanning based on new observations, and task completion checking, enhancing the ability to handle unexpected state changes of clothes (Section 5.2).
  • 3.
    Expansion of real-world experimental evaluations to diverse clothes categories, demonstrating the generality and robustness of our framework (Sec 6.4).
A.1 Open-category Semantic Keypoint Extraction
In the conference version, we trained a model to extract semantic keypoints from observations. However, due to the high-dimensional configuration space of clothes, collecting sufficient data is challenging. To mitigate data scarcity, we imported 3D clothes meshes into the simulation for data augmentation. Nevertheless, the limited availability of 3D mesh models restricted the model to extracting semantic keypoints from only four types of clothes. In this manuscript, we propose a novel open-category semantic keypoint extraction pipeline (Sec. 4) that leverages large vision-language models and foundation vision models to overcome data scarcity and eliminate the need for training on labeled data. In this new pipeline, introducing a new clothes category requires only a prototype image—without any annotations—for autonomous semantic keypoint discovery (Sec. 4.1). This enhancement significantly improves the applicability and generalization of the proposed method.
A.2 Closed-loop Task Planning
In the conference version, task planning was performed in symbolic space—without conditioning on visual observations—and executed only once at the beginning of the manipulation task. This made the system sensitive to unexpected state changes, which are common when the initial clothes state is highly crumpled, as in flattening tasks. In this manuscript, we propose a closed-loop task planning pipeline (Sec. 5.2), where the task planning is conditioned on visual observations. After each sub-task is executed, the observation is updated, and the pipeline determines whether to replan. Additionally, each generated plan is verified before execution, and the system continuously checks whether the overall task has been completed. If the task is not completed, a new plan is generated based on the current observation. This closed-loop task planning pipeline enhances the system’s ability to handle unexpected state changes of clothes. Experimental results in Sec. 6.2 show that this enhancement significantly improves the system’s performance when the initial clothes state is crumpled.
A.3 Expanded Real-world Experimental Evaluation
Beyond the qualitative real-world results on four clothing categories presented in the conference version, we have extended our real-world experiments to include quantitative evaluations and analysis across a broader range of clothes categories, varying in size, shape, and material (Sec. 6.4). These new results further demonstrate the effectiveness of our proposed clothes manipulation method, which demands a higher level of generalization and robustness.
By introducing a novel open-category semantic keypoint extraction pipeline and a closed-loop task planning system, this manuscript presents substantial extensions over the conference paper.
Appendix B Prompts
Basic Skill Discovery Define basic skills for a dual-arm home-service robot, which is required to perform clothes manipulation tasks like folding, flattening, hanging, and placing.
Think about this step by step: 1. provides some examples of clothes manipulation tasks. 2. breaks down these tasks into basic actions. 3. summarizes basic actions and derive a set of basic skills.
Semantic Keypoint Discovery Select the ’semantic keypoints’ of clothes from ’candidate keypoints’ to solve the clothes manipulation tasks. Don’t ignore any semantic keypoints.
The input request is an image of the clothes, annotated with a set of visual marks: candidate keypoints: black dots marked as ’P[i]’ on the image, where [i] is an integer.
The definition of semantic keypoints: keypoints that are crucial for clothes manipulation tasks like folding, flattening, and hanging, these keypoints can be described with natural language. Please note: In the folding process, both picking and placing keypoints are ’semantic keypoints’. Don’t ignore any semantic keypoints. Don’t ignore the center if the center is the placing point. clothes often have symmetric structures, which can be useful guidance.
The response should be a list of ’semantic keypoints’, each semantic keypoint should be a dictionary that contains: keypoint: P1 language description: ‘left shoulder’ reason: ‘crucial for symmetry and folding the top part along with the left shoulder’
Coarse Region Matching Match the region in the test image that corresponds to the specified keypoint in the template image.
The input request contains: template image annotated with a keypoint. The keypoint is marked as a red dot. test image that is divided into several regions annotated with a set of visual marks.
The response should be the mark of the corresponding region in the following format:
The region corresponding to the red keypoint is ** **.
Sub-task Proposal Describe the dual-arm robot’s actions to solve the clothes manipulation tasks by: selecting contact points from annotated semantic keypoints. selecting basic skills from a pre-defined skill library.
You have the following basic skills to select: grasp(): pick up clothes parts using the grippers of one or both arms, the parameter can be one or two contact points. moveto(): transports the clothes or clothes parts to a target position, the parameter can be one or two contact points. release(): drops clothes parts by opening grippers. rotate(): rotates the clothes grasped by the gripper. pull(): stretches the clothes to flatten them.
The input request contains: instruction: a task described in natural language. image: a top-down photo of clothes, annotated with semantic keypoints labeled as P[i] (where i is the index). semantic keypoint dictionary: a dictionary mapping each P[i] to its language description.
You should determine whether the task is completed and the sequence of subtasks based on annotated keypoints. Each subtask includes the selected basic skill and contact points.
The response should be a dictionary in JSON form as follows:
{ “completed”: false, “subtasks”: [ “grasp(left sleeve)”, “moveto(center)”, “release()”, “grasp(right sleeve)”, “moveto(center)”, “release()”, “grasp(left shoulder, right shoulder)”, “moveto(left hem, right hem)”, “release()” ] }
I will provide some examples to demonstrate how to use these basic skills and the response format.
Appendix C Additional Clothes Details
This section presents the parameters of the clothes used in our experiment. The clothes cover various categories, shapes, textures, sizes, and materials. Each clothes is assigned a unique ID, and its size and material are also listed in the table. The size information indicates the width and height of the clothes when fully flat. In addition, we capture an RGB image of each clothes from a top-down view.
No. 1
Size 47 × 40 cm
Material
Cotton & Polyester
No. 2
Size 55 × 50 cm
Material
Polyester
No. 3
Size 60 × 43 cm
Material
Cotton & Polyester
No. 4
Size 40 × 30 cm
Material
Cotton
No. 5
Size 112 × 58 cm
Material
Wool
No. 6
Size 53 × 42 cm
Material
Polyester & Spandex
No. 7
Size 46 × 35 cm
Material
Cotton & Polyester
No. 8
Size 53 × 41 cm
Material
Cotton & Spandex
No. 9
Size 18 × 30 cm
Material
Cotton
No. 10
Size 53 × 41 cm
Material
Nylon
No. 11
Size 72 × 51 cm
Material
Polyester & Spandex
No. 12
Size 46 × 57 cm
Material
Cotton & Spandex
No. 13
Size 70 × 40 cm
Material
Polyester
No. 14
Size 101 × 42 cm
Material
Silk
No. 15
Size 98 × 35 cm
Material
Viscose