### v2

# CLASP: General-Purpose Clothes Manipulation with Semantic Keypoints

Abstract
Clothes manipulation, such as folding or hanging, is a critical capability for home service robots. Despite recent advances, most existing methods remain limited to specific clothes types and tasks, due to the complex, high-dimensional geometry of clothes. This paper presents CLothes mAnipulation with Semantic keyPoints (CLASP), which aims at general-purpose clothes manipulation over diverse clothes types, T-shirts, shorts, skirts, long dresses, … , as well as different tasks, folding, flattening, hanging, … . The core idea of CLASP is semantic keypoints—e.g., “left sleeve” and “right shoulder”—a sparse spatial-semantic representation, salient for both perception and action. Semantic keypoints of clothes can be reliably extracted from RGB-D images and provide an effective representation for a wide range of clothes manipulation policies. CLASP uses semantic keypoints as an intermediate representation to connect high-level task planning and low-level action execution. At the high level, it exploits vision language models (VLMs) to predict task plans over the semantic keypoints. At the low level, it executes the plans with the help of a set of pre-built manipulation skills conditioned on the keypoints. Extensive simulation experiments show that CLASP outperforms state-of-the-art baseline methods on multiple tasks across diverse clothes types, demonstrating strong performance and generalization. Further experiments with a Franka dual-arm system on four distinct tasks—folding, flattening, hanging, and placing—confirm CLASP’s performance on real-life clothes manipulation. The CLASP source code is available online at https://github.com/dengyh16code/CLASP.
keywords:
Deformable object manipulation, state representation, foundation models
1 Introduction
People have long anticipated intelligent service robots performing home laundry chores: fold or hang shirts, pants, gowns … . While there has been significant progress in robot clothes manipulation in recent years (Avigal et al., 2022; Huang et al., 2022; Canberk et al., 2023; Longhini et al., 2024b), existing methods are usually tailored to specific clothes and tasks. General-purpose clothes manipulation—a single robot system performing different manipulation tasks over diverse clothes types—remains an important open challenge.
A core issue here is representation. Clothes are deformable objects with high-dimensional state space and complex geometry that varies significantly across categories, e.g., shirts and gowns. How do we represent clothes? Dense physics-based representations, such as particles (Bender et al., 2015; Chen et al., 2022), can model diverse clothes, but the fine details required are difficult to estimate from visual observations, because of clothes deformation and self-occlusion (Yin et al., 2021). Sparse keypoints (Gustavsson et al., 2022; Lips et al., 2024) are more robust to deformation, but generalize poorly over different clothes (Longhini et al., 2024a). Both approaches focus exclusively on geometry, while ignoring semantics. Semantics is crucial for clothes representation, as clothes are man-made objects designed specifically for humans.
We introduce semantic keypoints (Figure 1), e.g., “left sleeve” and “right shoulder”, as a spatial–semantic representation that connects perception and action for general-purpose clothes manipulation. Semantic keypoints provide significant advantages. First, they are salient for both perception and action. They can be extracted reliably with RGB-D sensors under moderate deformation and occlusion. They also identify commonly manipulated regions on clothes and highlight their affordance, providing an abstraction to guide action planning and execution. Second, semantic keypoint descriptors naturally emerge in human language due to their salience. Finally, semantic keypoints are grounded to structural features consistent within the same clothes category, enabling effective generalization.
Figure 1: CLothes mAnipulation with Semantic keyPoints (CLASP). The semantic keypoint representation enables CLASP to generalize over many different types of clothes and tasks. (a) Semantic keypoints for various types of clothes. (b) Four distinct clothes manipulation tasks in our experiments.
Semantic keypoint extraction poses two main challenges: (i) identifying semantically meaningful keypoints for open-category clothes, and (ii) achieving accurate spatial localization of keypoints under deformation and occlusion. To decouple the semantic and spatial challenges, we propose a two-stage method. First, we use a vision-language model (VLM) for semantic understanding of individual clothes categories. For each category, we determine the semantic keypoints on a prototype clothes instance lying flat. Next, we match these keypoints spatially to novel input clothes instances under possible deformation and occlusion, using vision foundation models.
Given a natural language instruction and RGB-D observations, CLothes mAnipulation with Semantic keyPoints (CLASP) performs various manipulation tasks over a wide range of clothes types. To tackle the challenge of general-purpose clothes manipulation, CLASP leverages semantic keypoints to decompose the manipulation task into a two-level hierarchy (Figure 2). At the high level, CLASP uses a VLM to transform the input language instruction into a task plan over the semantic keypoints. For example, grasp the left sleeve, move to the center, … . It verifies the generated task plan for geometric feasibility through motion planning and decides whether to replan. At the low level, CLASP executes each subtask in the high-level plan, using a set of pre-built manipulation skills conditioned on the keypoints. After executing each subtask, CLASP closes the loop by updating the visual observation and replanning if needed. The process continues until the task is successfully completed. To evaluate CLASP, we conducted experiments both in simulation and on a dual-arm robot system. For simulation experiments, we extended the SoftGym (Lin et al., 2021) benchmark and experimented with four clothes categories and four representative tasks under various scenarios. CLASP outperforms well-established baselines in clothes manipulation, including FlingBot (Ha and Song, 2022) and FabricFlownet (Weng et al., 2022). For real-robot experiments, we implemented CLASP on Franka robots for dual-arm manipulation on 15 clothes of various types. The results confirm the performance of CLASP and its generality. They demonstrate that semantic keypoints provide an effective representation of clothes manipulation, enabling general-purpose clothes manipulation with foundation models.
This work extends our earlier paper (Deng and Hsu, 2025) in three main aspects:
  • Semantic keypoint extraction for open-category clothes. We improved semantic keypoint extraction by leveraging foundation models, eliminating the need for labeled training data (Section 4). This significantly enhances the generalization performance of CLASP on open-category clothes.
  • Closed-loop planning. CLASP now closes the loop at both the task planning and the subtask execution levels (Section 5.2). Closed-loop planning substantially improves overall system performance (Sec. 6.2), especially with crumpled clothes.
  • Expanded real-world evaluation. We added diverse clothes, such as T-shirts, sweaters, shorts, pants, skirts, and dresses, and conducted comprehensive real-robot experiments to validate the effectiveness and generalization of CLASP (Section 6.4).
The rest of this paper is organized as follows. The related work is reviewed in Section 2. Section 3 introduces the overall CLASP framework. Section 4 presents semantic keypoints extraction. Section 5 details clothes manipulation guided by semantic keypoints. The experimental setup and results are provided in Section 6. Section 7 discusses limitations and future directions. Finally, Section 8 concludes the paper. Additional implementation and experimental details are provided in the appendix.
2 Related Work
2.1 State Representation for Deformable Object
Effective state representation is crucial for object manipulation. However, representing the state of deformable objects, such as clothes, remains a significant challenge due to their high-dimensional state space (Yin et al., 2021; Longhini et al., 2024a). Some task-specific state representations have been developed, such as using an approximate polygonal representation for clothes folding (Miller et al., 2012) and topological coordinates for assistive dressing (Wang et al., 2023). However, designing a general representation across different objects and tasks remains a significant challenge.
Physics-based models such as particle (Bender et al., 2015; Chen et al., 2022; Zhang et al., 2024b) and mesh (Guler et al., 2015; Macklin et al., 2016) provide a solution to model deformable dynamics and represent a wide range of objects. However, these dense models make accurate state estimation of deformable objects challenging, especially under deformation and occlusion (Lin et al., 2022).
Recently, learning to encode raw observations into a compact latent representation has emerged as a more effective solution (Lippi et al., 2020; Yan et al., 2021). However, they often struggle to generalize to new environments due to distribution shifts caused by variations in visual conditions (Hoque et al., 2022).
Keypoint representation provides a succinct way for deformable object representation, enabling efficient planning and learning (Kulkarni et al., 2019; Ma et al., 2022; Wu et al., 2024) while remaining robust to visual condition variations. Previous keypoints representation focuses on geometric information, using dense keypoints to represent the topology of deformable objects (Shi et al., 2021; Gustavsson et al., 2022; Zhou et al., 2023a). However, keypoint extraction is still challenged by occlusion and deformation, while dense keypoints often lack consistency and generalization. More recently, Huang et al. (2024) utilize vision foundation models to extract sparse keypoints and employs a vision-language model to impose keypoint-centric constraints for clothes manipulation, though it requires domain-specific human intervention.
In this paper, we present semantic keypoints as a general spatial-semantic representation of clothes. Semantic keypoints carry semantic meaning and can be described in natural language, making them sparse and salient for perception and action. Semantic keypoints are easy to extract, generalize well, and indicate where clothes are more likely to be manipulated.
2.2 Deformable Object Manipulation
Existing deformable manipulation approaches can be categorized into model-based and model-free approaches. Model-based approaches use a dynamic model to predict configurations of deformable objects under a certain action and then select the appropriate action (Cusumano-Towner et al., 2011; Miller et al., 2012; Yu et al., 2023). The key challenge lies in the dynamic modeling of deformable objects. To tackle this, two main techniques are used: physics-based models (Li et al., 2018; Chen et al., 2024), which rely on physical law, and data-driven models (Tanaka et al., 2018; Lin et al., 2022), which are learned from interaction data. Typical physics-based models include particle-based models (Longhini et al., 2024b; Zhang et al., 2024b) and mesh-based models (Huang et al., 2022). Recently, Material Point Method (MPM) (Jiang et al., 2016) has emerged as a new physics-based modeling method, which combines particles with grid-based forces for more accurate dynamic modeling (Gu et al., 2023). However, physics-based modes often struggle with heavy parameter tuning (Zheng et al., 2024). Data-driven dynamics models are learned from interaction data, typically formulated as a supervised learning problem (Hu et al., 2018; Tanaka et al., 2018; Yan et al., 2021). Compared to physics-based models, data-driven models support a more flexible state space, especially via latent representation (Lippi et al., 2020; Li et al., 2024). However, data-driven models rely on expensive robot interaction data and often fail to generalize to different objects and tasks. Overall, the generalization of model-based approaches highly depends on accurate dynamic modeling and calibration for unseen objects and tasks (Longhini et al., 2024a), which remains an open problem due to the complex, nonlinear dynamics of deformable objects.
Model-free approaches aim to learn robot actions directly without establishing a dynamic model, which can be categorized as imitation learning and reinforcement learning methods. The crucial challenge of imitation learning methods is collecting expert demonstrations, which can be gathered from human (Joshi et al., 2019; Lee et al., 2024) or scripted policies (Jia et al., 2019; Weng et al., 2022; Mo et al., 2022). Specifically, Avigal et al. (2022) employ imitation learning to acquire quasi-static action primitives, such as pick-and-place and dragging, for bimanual cloth folding. Lee et al. (2024) propose a method to learn fabric smoothing and folding policies from human demonstration videos. To avoid collecting extensive expert demonstrations, reinforcement learning methods offer an alternative solution. Matas et al. (2018) develop a sim-to-real reinforcement learning method for towel hanging. Ha and Song (2022) use reinforcement learning to learn a dynamic fling policy for efficient cloth flattening.
Although model-free approaches have made significant progress in a wide range of tasks (Ganapathi et al., 2021; Xu et al., 2022; Bahety et al., 2023), the generalization remains a significant limitation. Most of the previous methods aim to learn task-specific policies (Salhotra et al., 2022; Weng et al., 2022; Gao et al., 2023). Although some goal-conditioned methods use goal images for multi-task learning (Seita et al., 2021; Chi et al., 2022), they often struggle to generalize to unseen goals. To enhance generalization, collecting more data (Xue et al., 2023) and domain randomization (Hietala et al., 2022; Wang et al., 2023) have been explored. However, model-free approaches still struggle to generalize to unseen tasks or object categories. In this paper, we aim to tackle general-purpose clothes manipulation across diverse clothes categories and task types. To this end, we propose CLASP, a framework that integrates semantic keypoint representations with foundation models. CLASP enables generalization to a wide range of clothes and manipulation tasks.
Figure 2: CLASP Overview. Given an RGB-D observation, CLASP extracts semantic keypoints. These keypoints, along with the RGB image and task instruction, are fed to a VLM to generate a task plan. Once the plan is verified for feasibility through motion planning, the subtasks are executed sequentially. After each subtask execution, CLASP updates the observation and decides whether to replan. This process repeats until the overall task is completed.
2.3 Language-Conditioned Manipulation
Language-conditioned manipulation is an emerging field that aims to enable seamless interaction between humans and robots by teaching robots to understand and execute instructions specified in natural language (Shridhar et al., 2022; Mees et al., 2022; Zhou et al., 2023b). The recent emergence of large foundation models that are pretrained on internet-scale data, such as LLMs and VLMs, has significantly advanced language-conditioned manipulation tasks. These models provide commonsense knowledge for reasoning (Singh et al., 2023; Du et al., 2024) and perception (Huang et al., 2023; Gao et al., 2024), significantly enhancing generalization capabilities. More recently, vision–language–action models (VLAs) (Kim et al., 2024; Black et al., 2025; Pertsch et al., 2025) fine-tune VLMs for robot control, aiming for generalist policies but relying on large-scale robotic data.
Beyond specifying tasks, natural language also serves as a powerful abstraction of affordance (Mees et al., 2023), capturing the functions of an object, i.e., how it can be manipulated. For task-oriented grasping (Liu et al., 2020), GraspGPT (Tang et al., 2023) utilizes object class descriptions generated by a large language model (LLM) to identify optimal grasping regions. Similarly,  Ren et al. (2023) provide natural language descriptions of tools to accelerate learning language-conditioned tool manipulation policies. RoboPoint finetunes a VLM for spatial affordance prediction conditioned on natural language (Yuan et al., 2024). MimicFunc (Tang et al., 2025) leverages functional cues to identify keypoints for tool manipulation.
Most previous language-conditioned object manipulation methods focus on rigid objects, with few addressing deformable object manipulation (Deng et al., 2024; Barbany et al., 2025). For rigid objects, existing methods are available to execute actions such as picking, placing, moving, and grasping (Liang et al., 2023). However, a significant gap remains between task planning and action execution when it comes to deformable object manipulation. Motivated by prior work that uses natural language as an abstraction of affordance, we propose semantic keypoints as a spatial-semantic representation of clothes. These keypoints act as an interface between VLM-powered task planning and low-level action execution, enabling general-purpose clothes manipulation.
3 Overview
For general-purpose clothes manipulation, CLASP leverages semantic keypoints to decompose the task into a two-level hierarchy. Figure 2 illustrates the overall framework. Given RGB-D observations of clothes and a language instruction, CLASP generates a sequence of manipulation actions to complete the specified task.
First, CLASP extracts semantic keypoints using foundation models. Each extracted keypoint is represented by a 3D position and an associated language descriptor. These keypoints, together with the RGB image and task instruction, are then fed into a vision-language model (VLM) to generate a task plan over semantic keypoints (e.g., grasp(left sleeve), moveto(center), …). Each subtask in the plan is grounded by selecting the appropriate skill from a pre-built skill library. The plan is then verified for feasibility through motion planning, which generates the corresponding robot trajectories for execution. After each subtask, CLASP updates its observation and determines whether replanning is required. This closed-loop process continues until the task is successfully completed.
In the following sections, we will introduce how we extract effective semantic keypoints of open-category clothes in Sec. 4. Then, to enable general-purpose clothes manipulation with semantic keypoints, we present the construction of the skill library in Sec. 5.1, followed by the details of task planning and execution in Sec 5.2.
Figure 3: Semantic keypoint discovery. A fully-automated pipeline discovers semantic keypoints on a prototype image for each clothes type.
4 Semantic Keypoints
The core idea of CLASP is to use semantic keypoints as a general state representation for clothes. Each semantic keypoint includes: (i) a language description (e.g., “left sleeve”), providing semantic context for high-level task planning, and (ii) a keypoint position, offering geometric information to guide low-level action generation. Although previous work has explored using keypoints for clothes manipulation, most methods rely on manually annotated keypoints (Mo et al., 2022; Lips et al., 2024). Since clothes are highly deformable, collecting sufficient data that captures clothes’ various configurations is time-consuming and labor-intensive. To mitigate data scarcity, augmenting 3D synthetic data in the simulation environment has been explored (Deng and Hsu, 2025). However, this approach is difficult to scale to open-category clothes due to the limited availability of 3D clothes models. Recently, researchers have explored zero-shot keypoint extraction by leveraging powerful and generalizable feature maps from vision foundation models (Huang et al., 2024; Di Palo and Johns, 2024; Fang et al., 2024b). However, these methods can not ensure that the resulting keypoints have explicit semantic meaning and can be described using natural language. In addition, using the vision foundation model for consistent keypoint extraction on clothes remains challenging, especially under self-occlusion and deformation.
Given that semantic keypoints encode both semantic and spatial information, two key technical challenges must be addressed in keypoint extraction: (1) understanding the semantics of open-category clothes and (2) maintaining the spatial precision of keypoints under deformation and occlusion. To address them, we propose a two-stage pipeline for keypoint extraction. The first stage (Figure 3) focuses on semantic understanding, aiming to autonomously discover semantic keypoints on prototype images and describe them using natural language. Specifically, we leverage geometric prior and the vision language model’s commonsense knowledge to develop an autonomous semantic keypoint discovery pipeline. For each clothes category, we provide a prototype image free from occlusion and deformation, which makes semantic understanding feasible and also provides full observation as a reference in the second stage. The second stage (Figure 4) focuses on spatial precision, aiming to match semantic keypoints on the prototype image to novel clothes, which may involve occlusions and deformations. Specifically, we utilize a coarse-to-fine matching strategy to propose semantic keypoints on novel clothes precisely.
Figure 4: Semantic keypoint matching. Given an image of novel clothes, we first retrieve the most relevant prototype. Each semantic keypoint on the prototype is then matched to the novel clothes through a coarse-to-fine pipeline. Specifically, a VLM is employed for coarse region matching, while Stable Diffusion (SD) and DINOv2 are utilized for fine-grained keypoint matching.
4.1 Keypoint Discovery
For each clothes type, we provide a prototype image free from occlusion and deformation, allowing the semantic keypoint discovery stage to focus on understanding semantics and defining semantic keypoints for each clothes type. To discover semantic keypoints on the prototype image, we design an autonomous pipeline as shown in Figure 3, which leverages the geometric prior of clothes and the semantic understanding capability of the VLM for keypoint discovery.
To discover semantic keypoints, we first extract the contour and skeleton of clothes, which serve as an effective abstraction, capturing the geometric structure of clothes and encoding the human body. Specifically, we employ OWL-V2 (Minderer et al., 2024) to generate the bounding box of clothes on the prototype image based on the text prompt. Later, the bounding box will serve as the box prompts for the Segment-Anything Model (SAM) (Kirillov et al., 2023) to generate the corresponding masked image $I_{m}$ of the clothes. We then extract the contour of clothes from the masked image $I_{m}$ and perform skeletonization by calculating equidistance to the boundaries (Lee et al., 1994), resulting in the skeleton. To get keypoints from the skeleton and boundary, we then apply farthest point sampling (fps) (Eldar et al., 1997) on the boundary and skeleton to sample candidate keypoints $\boldsymbol{K}_{c}$. Farthest point sampling is an effective method for obtaining diverse, spatially distributed candidate keypoints that are geometrically salient.
After obtaining candidate keypoints $\boldsymbol{K}_{c}$, selecting semantic keypoints requires semantic understanding of clothes and clothes manipulation. Therefore, we leverage the commonsense knowledge of the VLM and employ mark-based visual prompting (Fang et al., 2024a) to select semantic keypoints and describe them with natural language. In this paper, we use GPT-4o as the VLM. The prompt is multi-modal and consists of two parts: (i) text prompt, and (ii) marked image. In the text prompt, we provide context information to the VLM, including the definition of semantic keypoints, the expected behavior of the VLM, and the desired response format. In the marked image, we draw visual marks on candidate keypoints $\boldsymbol{K}_{c}$ and label them with numbers. Through mark-based visual prompting, we can get the prototype semantic keypoints $\boldsymbol{K}_{0}=\{p_{i,0},\,s_{i,0}\}_{i=1}^{n}$, where $p_{i,0}$, $s_{i,0}$ denote the pixel position and language description of $i$-th semantic keypoint, respectively, and $n$ is the number of keypoints.
In this way, we establish a prototype set, where each image represents a clothes type and is annotated with semantic keypoints.
4.2 Keypoint Matching
Although the autonomous pipeline can discover semantic keypoints on the prototype image in the first stage, it does not perform robustly on clothes with occlusion and deformation, which may lead to failures in boundary and skeleton extraction. Thus, the second stage performs semantic keypoint matching to transfer the semantic keypoints from the prototype image to novel clothes. The second stage ensures precise semantic keypoints extraction under occlusion and deformation.
The pipeline of semantic keypoint matching is shown in Figure 4. Given an RGB observation image of novel clothes, we first retrieve the most relevant prototype image by computing the Best-Buddies Similarity (BBS) using features extracted from the vision foundation model DINOv2 (Oquab et al., 2023). DINOv2 can extract rich and transferable visual features efficiently. BBS is a robust metric for measuring structural similarity between two images. It identifies pairs of local patches—one from the observation image and one from the prototype image—that are mutual nearest neighbors in the feature space. These pairs, referred to as best buddy pairs, capture strong local correspondences. The BBS score is defined as the proportion of local patches in the observation image that form such pairs. By evaluating the BBS score across all prototype images, we retrieve the one that is most structurally similar to the observation image. To enhance robustness, we apply a set of rotations to the observation image and select the rotation that yields the highest BBS score.
If the highest BBS score within the prototype image set falls below a predefined threshold, the observation is classified as a new clothes type, requiring a new prototype image. This new prototype image is then annotated using our autonomous semantic keypoint discovery pipeline. Otherwise, the prototype image with the highest BBS score is retrieved.
After retrieving the most relevant prototype image $I_{0}$ annotated with semantic keypoints, we then perform a coarse-to-fine semantic keypoint matching on novel clothes. Given the novel clothes’ RGB observation image $I$ and a semantic keypoint position $p_{i,0}$ on the prototype image, we first match $p_{i,0}$ to a coarse region $\Omega$ on the observation image $I$. This stage is crucial for capturing global context and avoiding local minima. Specifically, we employ VLM with mark-based visual prompting for coarse region proposals. We first mark the observation image $I$ with an overlaid grid, where each grid cell is indexed by a unique text label, and mark the prototype image $I_{0}$ with a dot on $p_{i,0}$. We then prompt the VLM with the marked observation and prototype image to select the grid indices corresponding to the prototype semantic keypoint position $p_{i,0}$. To obtain precise semantic keypoint positions, we then perform fine-grained keypoint matching on the region $\Omega$ with vision foundation models. Borrowing the method from (Zhang et al., 2024a), we fuse the feature maps from DINOv2 (Oquab et al., 2023) and Stable Diffusion (SD) (Rombach et al., 2022). The fused feature effectively captures both local and global information, thereby enhancing the robustness of semantic keypoint matching. Specifically, we feed the observation image $I$ and the prototype image $I_{0}$ into the fused feature extractor $\mathcal{F}(\cdot)$, which produces their corresponding fused feature maps $\mathcal{F}(I)$ and $\mathcal{F}(I_{0})$. For each point $p\in\Omega$ on the observation image, we obtain its feature representation $\mathcal{F}(I,\,p)$, and for each prototype semantic keypoint $p_{i,0}$, we obtain its corresponding prototype feature $\mathcal{F}(I_{0},\,p_{i,0})$. We then compute the cosine similarity between $\mathcal{F}(I,\,p)$ and $\mathcal{F}(I_{0},\,p_{i,0})$ to perform fine-grained keypoint matching:
$p_{i}=\arg\max_{p\in\Omega}\langle\mathcal{F}(I,\,p),\,\mathcal{F}(I_{0},\,p_{i,0})\rangle,$ (1)
where $\langle\cdot,\cdot\rangle$ denotes the cosine similarity between two feature vectors. Here, $p_{i}$ represents the pixel position of the matched keypoint on the observation image $I$ corresponding to the prototype keypoint $p_{i,0}$. Repeating this process for all prototype semantic keypoints $\boldsymbol{K}_{0}$, we obtain the matched semantic keypoints $\boldsymbol{K}$ on the novel clothes. With the pixel locations of these keypoints, we further compute their 3D coordinates using the corresponding depth map and camera parameters.
Figure 5: Skill library. CLASP skill library consists of 5 basic skills: grasp, moveto, release, rotate and pull.
5 Clothes Manipulation With Semantic Keypoints
5.1 Skill Library
For general-purpose clothes manipulation, we build a skill library first. Instead of manually designing skills, we prompt an LLM to discover basic clothes manipulation skills. The LLM’s commonsense knowledge ensures these skills are general and sufficient for various clothes and manipulation tasks. Specifically, the LLM is prompted with a chain-of-thought strategy (Wei et al., 2022) to (i) generate examples of clothes manipulation tasks, (ii) break down these tasks into basic actions, and (iii) summarize the actions identified in step (iii) to derive a set of basic skills. In this way, we identify the basic skill set $\mathcal{A}$, including grasp: pick up a part of the clothes, moveto: transport the clothes to a target position, release: drop a part of the clothes, rotate: rotate the clothes, and pull: stretch the clothes for flattening. Figure 5 visualizes the basic skills discovered by the LLM. Since semantic keypoints serve as the effective affordances for clothes manipulation and define where clothes are more likely to be manipulated, we implement these basic skills as policies parameterized by contact point positions. Specifically, we define heuristic rules to create waypoints based on the contact point positions for each basic skill. Once the waypoints are determined, a motion planning algorithm is applied to generate the complete trajectory for the robot’s execution. Specifically, we implement each basic skill as follows:
[…]
release()(): The skill release drops clothes parts by opening grippers. After the release action, the observation is updated, and CLASP extracts new semantic keypoints.
[…]
pull(l)(l): The skill pull stretches the clothes to flatten them, where $l$ represents the stretch distance. In this paper, we set the stretch distance $l$ to 10% of the distance between the grippers of the two arms. A potential improvement could involve stretching the clothes as much as possible without tearing, monitored through a force sensor or a front-view camera.
5.2 Task Planning
Figure 6: Example of few-shot prompting for the VLM-powered task planner.
Algorithm 1 Task Planning Procedure
1:Instruction $\mathcal{L}$, image $I$, keypoints $\boldsymbol{K}$, skill set $\mathcal{A}$
2:Valid task plan or failure reason
3:if CheckCompletion($\mathcal{L},I,\boldsymbol{K}$) then
4:  return Task Completed
5:end if
6:repeat
7:  $\boldsymbol{W}\leftarrow$ ProposeTaskPlan($\mathcal{L},I,\boldsymbol{K},\mathcal{A}$)
8:  $valid\leftarrow$ PlanVerification($\boldsymbol{W},\mathcal{A},\boldsymbol{K}$)
9:until validvalid or attempts are exhausted
10:return $\boldsymbol{W}$ or failure reason
Algorithm 2 Check validity of a proposed task plan $\boldsymbol{W}$
1:PlanVerification($\boldsymbol{W},\mathcal{A},\boldsymbol{K}$)
2:for each $\boldsymbol{w}_{j}=(a_{j},c_{j}^{1},c_{j}^{2})$ in $\boldsymbol{W}$ do
3:  if $a_{j}\notin\mathcal{A}$ or $c_{j}^{1,2}\notin\boldsymbol{K}$ then
4:   return False
5:  else if not MotionFeasibility($a_{j},c_{j}^{1},c_{j}^{2}$) then
6:   return False
7:  end if
8:end for
9:return True
After establishing the skill library for clothes manipulation, we leverage the VLM to perform task planning over semantic keypoints, utilizing its commonsense knowledge acquired from internet-scale data (Zhao et al., 2024). The task planning procedure (Algorithm 1) consists of three main modules: CheckCompletion, ProposeTaskPlan, and PlanVerification.
CheckCompletion: Given a language instruction $\mathcal{L}$, an observation image $I$, and extracted semantic keypoints $\boldsymbol{K}$, CLASP first determines whether the task specified by $\mathcal{L}$ has already been completed. If the completion condition is satisfied, the procedure terminates.
ProposeTaskPlan: If the task remains incomplete, the VLM generates a candidate task plan $\boldsymbol{W}=(\boldsymbol{w}_{1},\boldsymbol{w}_{2},\ldots,\boldsymbol{w}_{H})$. Each subtask $\boldsymbol{w}_{j}=(a_{j},c_{j}^{1},c_{j}^{2})$ consists of a basic skill $a_{j}$ and up to two contact points $(c_{j}^{1},c_{j}^{2})$ for the dual-arm manipulation system. The skill $a_{j}$ is selected from the predefined set $\mathcal{A}$ introduced in Sec. 5.1, while the contact points are chosen from the semantic keypoints $\boldsymbol{K}$. Semantic keypoints define where clothes can be manipulated and serve as contact point candidates in high-level task planning. To improve robustness, we employ in-context learning by providing the VLM with a few-shot prompt that includes example pairs of natural-language instructions, observation images annotated with semantic keypoints, and corresponding task plans (Figure 6). These examples define the expected input–output structure and demonstrate how to ground instructions into executable skills over semantic keypoints.
PlanVerification: The proposed candidate task plan $\boldsymbol{W}$ is then validated through the procedure described in Algorithm 2. For each subtask $\boldsymbol{w}_{j}$, CLASP first performs syntax and semantic checks—verifying that the skill $a_{j}$ exists in $\mathcal{A}$ and that contact points $c_{j}^{1,2}$ correspond to valid semantic keypoints. Next, it evaluates geometric feasibility via motion planning. A subtask passes verification if a feasible trajectory is found; otherwise, the failure reason is used to prompt VLM to generate a new task plan.
Once a verified task plan is obtained, CLASP executes the subtasks sequentially. After each subtask, the RGB-D observation is updated. However, due to potential occlusion caused by the robot, semantic keypoints are re-extracted only after executing a release action. Following each keypoint update, the VLM is prompted to replan based on the updated observation and keypoints. This process repeats iteratively until the task is successfully completed.
6 Experiments
Figure 7: Qualitative results on semantic keypoint extraction. Our method can achieve open-category semantic keypoint extraction under irregular deformation and occlusion.
Our experiments investigate three questions:
Q1. Can CLASP’s submodules perform effectively under varying conditions?
We evaluate the performance of CLASP’s submodules: semantic keypoint extraction and task planning. Sec. 6.1 shows that the proposed method reliably extracts semantic keypoints across various clothes, even under occlusion and deformation. Sec. 6.2 shows that the task planner generates robust, executable plans, with closed-loop planning substantially improving performance on crumpled clothes.
Q2. How well does CLASP perform on clothes manipulation compared with the existing baseline method?
We address this question through comparative experiments detailed in Sec. 6.3. CLASP outperforms baseline methods in terms of manipulation success rate and generalization capability, highlighting its effectiveness.
Q3. Can CLASP generalize to a wide range of clothes types and manipulation tasks in the real world?
To answer this, we conduct extensive real-world evaluations, as presented in Sec. 6.4. Results show that CLASP generalizes well to various clothes types and manipulation tasks, confirming its practical applicability.
6.1 Semantic Keypoint Extraction Experiments
Table 1: Semantic keypoint extraction performance.
Expert AKD (pixel) $\textit{AP}_{15}$ (%) $\textit{AP}_{30}$ (%) $\textit{AP}_{45}$ (%)
1 16.2 71.0 88.0 90.7
2 17.8 69.0 84.1 90.4
3 19.7 68.7 82.2 87.8
We first evaluated the performance of CLASP in extracting effective semantic keypoints. Since no ground-truth semantic keypoints are available, we provide annotated prototype images to three human experts and ask them to label semantic keypoints on images of novel clothes. In this way, we establish a testing set of 120 images from 15 different clothes. In these images, clothes have diverse configurations, poses, occlusions, and deformations. We then compare the semantic keypoints extracted by CLASP against those labeled by the experts.
The evaluation utilizes two standard metrics for keypoint extraction: Average Keypoint Distance (AKD) and Average Precision (AP). AKD calculates the average distance between the extracted keypoints and the ground-truth keypoints, while AP measures the proportion of correctly extracted keypoints under different thresholds. Since the resolution of the observed RGB-D images is 720×1280, we used thresholds of 15, 30, and 45 pixels for AP calculation.
The results are shown in TABLE 1. The average keypoint distance between the semantic keypoints proposed by human experts and our method is less than 20 pixels, while the average precision under the 30 pixels threshold is more than 80%.
Considering the resolution and uncertainty in human expert labeling, such as identifying the center of clothes, the results can demonstrate that our proposed method performs well in extracting semantic keypoints. Figure 7 shows some qualitative results of semantic keypoint extraction. The experiment results demonstrate that our two-stage semantic keypoint extraction pipeline is effective in extracting semantic keypoints of clothes even under occlusion and deformation. The prior knowledge from the VLM and vision foundation models ensures the generalization to diverse clothes types. Besides, coarse-to-fine keypoint matching ensures spatial precision and makes semantic keypoint extraction robust. Our method enables effective semantic keypoint extraction on open-category clothes using a single unannotated prototype image per clothes category, without the need for large-scale labeled data.
Table 2: Task planning performance.
Task Check Completion (%) Subtask Accuracy (%)
VLM VLM + KP LLM VLM VLM + KP
Folding 66.7 93.3 33.3 56.7 86.7
Flattening 73.3 96.7 50.0 60.0 93.3
Hanging 93.3 100.0 96.7 76.7 100.0
Placing 83.3 96.7 96.7 70.0 96.7
6.2 Task Planning Experiments
To evaluate the performance of the proposed task planner for clothes manipulation, we collect images annotated with ground-truth semantic keypoints (labeled by human experts) along with free-form natural language descriptions to evaluate the planner on folding, flattening, hanging, and placing tasks. For each task category, we collect 30 images of 15 diverse clothes. We evaluate the planner using two metrics: check completion accuracy, which assesses whether the planner correctly determines if the overall task has been completed based on the observation; and subtask accuracy, which measures whether the generated subtask sequence is appropriate for the observed state. We compare our method (VLM + KP) with two baselines: (i) a VLM baseline that takes raw images without semantic keypoints, and (ii) an LLM baseline that receives only the language descriptors of semantic keypoints along with the language instruction, without any visual input (as in our previous work Deng and Hsu (2025)). All methods receive the same few-shot demonstrations but differ in visual input: our method uses keypoint-annotated images, the VLM baseline uses raw images, and the LLM baseline does not rely on any visual input.
Table 2 presents the experimental results. Since the LLM baseline lacks visual input, we do not report its accuracy in checking completion. It performs the worst on folding and flattening tasks, where adaptation to varied states is crucial. For folding, some parts of the clothes (such as sleeves) may be folded or unfolded after execution. For flattening, the model needs to understand the state of the clothes from visual observation and determine whether flattening is required. For example, in some flattening tasks, the robot must repeatedly flatten the clothes to achieve a flat state. The LLM baseline, being an open-loop planner, generates task plans solely from symbolic descriptions without observing the outcomes of action execution. This open-loop nature leads to failures in tasks that require dynamic replanning. In contrast, our planner operates in a closed-loop manner, continuously updating the plan based on visual and keypoint feedback to adjust its strategy when action outcomes deviate from expectations. The LLM baseline performs slightly better in hanging and placing tasks, which involve less uncertainty in the initial state. Moreover, the LLM’s symbolic reasoning remains consistent and unaffected by visual distribution shifts.
Compared to the VLM baseline without semantic keypoints, our pipeline achieves more reliable and interpretable task planning, especially in long-horizon folding tasks. Without semantic keypoints, the VLM often hallucinates or generates incorrect plans. Semantic keypoints provide critical structural information about the clothes’ state, allowing the model to reason over geometry and topology rather than surface appearance. This enhances robustness under distribution shifts between the observed images and the few-shot examples in the prompt.
Overall, our proposed planner achieves reliable accuracy in generating appropriate task plans and determining task completion. By providing semantic keypoints and a predefined skill library, we enable structured and controllable planning. To further enhance robustness, we employ in-context learning with examples that include language instructions, observation images with semantic keypoint annotations, and corresponding task plans. All generated plans are also verified before execution, and the planner supports dynamic replanning based on new observations, resulting in stable and robust closed-loop task planning.
Table 3: Simulation experiments. The average success rates (%) on testing tasks. The best performance is in bold.
Method
Folding
(seen object)
Flattening
(seen object)
Hanging
(seen object)
Placing
(seen object)
Towel T-shirt T-shirt Skirt Trousers Towel Towel Skirt
CLIPORT (Shridhar et al., 2022) 77.5 80.0 32.5 36.7 76.7 83.3 93.3 93.3
Goal-conditioned Transporter (Seita et al., 2021) 83.3 76.7 26.7 33.3 100.0 80.0 66.7 83.3
FlingBot (Ha and Song, 2022) N/A N/A 66.7 85.0 N/A N/A N/A N/A
FabricFlowNet (Weng et al., 2022) 93.7 100.0 N/A N/A N/A N/A N/A N/A
CLASP 100.0 95.0 65.0 80.0 96.7 97.5 96.7 96.7
Method
Folding
(unseen object)
Flattening
(unseen object)
Hanging
(unseen object)
Placing
(unseen object)
Trousers Skirt Trousers Towel T-shirt Skirt Trousers T-shirt
CLIPORT (Shridhar et al., 2022) 0.0 0.0 8.3 9.2 76.7 66.7 70.0 73.3
Goal-conditioned Transporter (Seita et al., 2021) 0.0 0.0 10.0 6.7 36.7 40.0 36.7 60.0
FlingBot (Ha and Song, 2022) N/A N/A 29.2 34.2 N/A N/A N/A N/A
FabricFlowNet (Weng et al., 2022) 0.0 2.5 N/A N/A N/A N/A N/A N/A
CLASP 87.5 81.7 60.8 65.0 93.3 76.7 93.3 94.2
6.3 Simulation Experiments
To evaluate CLASP’s performance on clothes manipulation, we conduct experiments in SoftGym, where clothes are modeled as particles with ground-truth positions. We sample 3D meshes of clothes from CLOTH3D (Bertiche et al., 2020) dataset, covering four common clothes types in human’s daily life: T-shirts, trousers, skirts, and towel. For each type, over 35 instances of varying sizes and shapes are used. In addition, we extend the SoftGym benchmark to 16 clothes tasks. These tasks can be divided into 4 types:
  • Folding tasks involve folding clothes to achieve a goal configuration. The success of a folding task is determined by the mean particle position error between the achieved and goal configuration.
  • Flattening tasks involve flattening crumpled clothes with random deformations and occlusion. The success of a flattening task is determined by the coverage area of the clothes.
  • Hanging tasks require hanging clothes on a rack. A hanging task is successful when the clothes are fully hung on the rack without any part touching the ground.
  • Placing tasks require placing the clothes in a box. A placing task is successful when the clothes are laid flat inside the box.
Baseline methods include two general multi-task learning frameworks (CLIPORT and Goal-conditioned Transporter) and two task-specific algorithms (FlingBot and FabricFlowNet):
  • CLIPORT (Shridhar et al., 2022) represents a typical end-to-end algorithm for learning language-conditioned manipulation policies, which leverages a pretrained vision language model and a two-stream architecture for semantic understanding and spatial precision.
  • Goal-conditioned Transporter (Seita et al., 2021) represents a typical goal-conditioned transporter network for deformable object manipulation, which infers the optimal action from the current and goal images with a TransporterNets structure.
  • FlingBot (Ha and Song, 2022) is a self-supervised learning framework designed for flattening crumpled clothes by using an effective fling action.
  • FabricFlowNet (Weng et al., 2022) is a learning framework designed for clothes folding, which infers the optimal action based on optical flow.
For each baseline method, we provide 1000 expert demonstrations per task for model training. To evaluate the generalization of baseline methods, only half of the tasks are seen during training through expert demonstrations. For CLASP, we provide few-shot examples of seen tasks in the prompt. Unseen tasks involve new object types. For each task, we conduct 120 trials with different clothes configurations to calculate the success rate. Two task-specific baseline methods are designed for clothes folding and flattening, respectively. We only test them on related tasks, and other tasks are marked as N/A (Not Applicable).
The experiment results are shown in TABLE 3. Overall, CLASP outperforms the two multi-task learning methods on both seen and unseen tasks. Compared with Goal-conditioned Transporter, CLIPORT shows better generalization. Language instructions facilitate capturing similarities between different tasks at the symbolic level. Besides, the pretrained vision-language model enables CLIPORT to capture such similarities. However, CLIPORT’s generalization is limited when adapting to long-horizon tasks, such as transferring the skill from folding a T-shirt to folding trousers, because it learns task-specific manipulation action sequences in an end-to-end manner, and the learned action sequences are not transferable. In contrast, CLASP learns task-agnostic and generalizable language and visual concepts across a wide range of clothes manipulation tasks. The commonsense knowledge from VLM allows CLASP to handle clothes manipulation tasks by decomposing them into predefined basic skills. Furthermore, semantic keypoints are a general spatial-semantic representation of clothes, which are task-agnostic and provide effective cues for task planning and action generation in clothes manipulation.
Compared to the two task-specific algorithms, CLASP shows comparable performance on seen clothes folding and flattening tasks, demonstrating the effectiveness of the proposed method for clothes manipulation. However, on unseen clothes types, two task-specific algorithms tend to fail due to distribution shifts in observations. In contrast, CLASP is a general-purpose clothes manipulation method.
Figure 8: Robot experiments. The experimental setup consists of a Franka dual-arm system performing folding, flattening, hanging, and placing over many different clothes types.
Table 4: Robot experiments. The success rates on various clothes and tasks.
Task Clothes ID
1 2 3 4 5 6 7 8 9 10 11 12 13 14 15 Mean
Folding 4/5 5/5 4/5 4/5 5/5 5/5 4/5 5/5 3/5 4/5 4/5 5/5 5/5 3/5 4/5 4.3/5
Flattening 2/5 4/5 4/5 3/5 3/5 4/5 2/5 5/5 2/5 3/5 3/5 4/5 3/5 3/5 4/5 3.3/5
Hanging 5/5 5/5 5/5 5/5 N/A 5/5 4/5 5/5 5/5 4/5 4/5 5/5 N/A N/A N/A 4.7/5
Placing 4/5 5/5 5/5 5/5 N/A 5/5 5/5 5/5 5/5 4/5 3/5 5/5 N/A N/A N/A 4.6/5
6.4 Real-world Experiments
To evaluate the performance and generalization of CLASP in real-world scenarios, we establish a dual-arm robot manipulation system. As shown in Figure 8, the system consists of two Franka Research 3 robot arms and a top-down RealSense L515 RGB-D camera for capturing RGB-D images. The clothes are placed on a table in front of the robot. We use SAM (Kirillov et al., 2023) and OWLv2 (Minderer et al., 2024) to segment the clothes for semantic keypoint extraction. To generate the dual-arm trajectories, we implement a motion planning algorithm using MoveIt! (Chitta et al., 2012) to avoid collisions and synchronize the motion between the arms.
In real-world experiments, we evaluate CLASP on 15 clothes across diverse types, such as shirts, T-shirts, sweaters, skirts, long dresses, and vests, size (ranging from 18 cm × 30 cm for baby pants to 42 cm × 101 cm for adult long dresses), shapes (long sleeves, short sleeves, and sleeveless), and materials (e.g., cotton, polyester, and wool). We directly evaluate CLASP’s performance on manipulation tasks of these clothes. Each clothes item is evaluated across four tasks: folding, flattening, hanging, and placing. The evaluation process includes five trials per task, with success rates calculated as follows. Folding: Success is determined by the Intersection over Union (IoU) between cloth masks achieved by CLASP and human experts. If the IoU is 90% or higher, the task is considered successful. Flattening: if the flattened clothes cover at least 90% of their maximum possible area in a fully flat state, the flattening task is successful. Hanging: the task is considered successful if the clothes are completely hung on the rack without any part touching the ground. Placing: the task is considered successful when the clothes are entirely placed within the box.
The experiment results are shown in Table 4. For clothes 5, 13, 14, and 15, due to limitations in the robot’s working space, we did not perform evaluations on hanging and placing tasks. These are marked as N/A (Not Applicable). Overall, CLASP can be applied to a wide variety of clothes and manipulation tasks while achieving a high success rate. CLASP achieves an 86% success rate in clothes folding, a 66% success rate in clothes flattening, a 94% success rate in clothes hanging, and a 92% success rate in clothes placing. The success rate is comparable to existing task-specific clothes manipulation algorithms, while we test CLASP on broader clothes types and instances. With a general semantic keypoints representation and the prior knowledge from foundation models, CLASP is an effective general-purpose clothes manipulation method. Fig. 9 illustrates some representative examples. For each task, we visualize the extracted semantic keypoints, manipulation action sequences, and achieved object state. CLASP demonstrates the ability to manipulate various clothes in different ways in real-world scenarios. More videos of real-world experiments can be found in our supplementary material.
Figure 9: Qualitative results of real-world experiments. From left to right, the figure displays the task description, extracted semantic keypoints, the real robot’s execution process, and the final achieved state.
7 Discussion
Despite the promising overall results, CLASP still shows several limitations based on the experimental results. First, CLASP struggles to extract accurate semantic keypoints of clothes with significant occlusion or deformation. This limitation arises because such cases are underrepresented in the datasets of the vision foundation models and VLMs that CLASP relies on. Although CLASP can perform flattening based on semantic keypoints to reduce occlusion, its success rate is lower than in other tasks (Table 3 and Table 4). Nevertheless, it achieves such performance and generalization without using any robot manipulation data. Future work could improve robustness by refining the policy through imitation learning with additional demonstrations. Second, CLASP faces challenges in manipulating large clothes due to workspace limitations. In real-world experiments, it sometimes fails to generate geometric feasible task plans for large clothes, preventing evaluation on certain hanging and placing tasks (e.g., clothes 5, 13, 14, and 15). This limitation could be addressed by integrating a mobile manipulator to expand the reachable workspace (Yang et al., 2024) or by adding dragging skills (Toussaint et al., 2018), allowing CLASP to handle larger clothes with random pose. Lastly, during manipulation, sliding or deformation can affect performance, especially for silk clothes (clothes 14 in real-world experiments). The folding success rate remains relatively low due to excessive sliding and deformation. This issue could be addressed by closed-loop action execution based on real-time semantic keypoint tracking. However, since CLASP’s keypoint extraction depends on large foundation models, real-time perception is still challenging. A practical improvement would be to use CLASP’s current pipeline to generate labeled data and distill a smaller, lightweight keypoint extractor, enabling real-time keypoint tracking and online trajectory optimization to enhance action robustness.
8 Conclusion
In this paper, we present semantic keypoints as a general spatial-semantic representation of clothes. Building on this representation, we propose CLothes mAnipulation with Semantic keyPoints (CLASP), where semantic keypoints play a crucial role in both high-level task planning and low-level action generation. By integrating the general semantic keypoint representation with foundation models pretrained on internet-scale data, CLASP serves as a general-purpose method for clothes manipulation. Simulation experiments demonstrate that our method outperforms baseline methods in terms of success rate and generalization in clothes manipulation. Real-world experiments further validate CLASP’s generalization in clothes manipulation. CLASP can be directly applied to a diverse range of clothes types and manipulation tasks in real-world scenarios.
Appendix A Discussion of Differences with the Conference Version
This manuscript is an extension of our previous work presented at the IEEE International Conference on Robotics and Automation (ICRA), titled “General-purpose Clothes Manipulation with Semantic Keypoints” (Deng and Hsu, 2025). We have significantly expanded upon the conference version in three primary aspects:
  • 1.
    Introduction of a new semantic keypoint extraction pipeline for open-category clothes, enhancing applicability and generalization (Section 4).
  • 2.
    Introduction of a closed-loop task planning pipeline capable of plan verification, dynamic replanning based on new observations, and task completion checking, enhancing the ability to handle unexpected state changes of clothes (Section 5.2).
  • 3.
    Expansion of real-world experimental evaluations to diverse clothes categories, demonstrating the generality and robustness of our framework (Sec 6.4).
A.1 Semantic Keypoint Extraction for Open-category Clothes
In the conference version, we trained a model to extract semantic keypoints from observations. However, due to the high-dimensional configuration space of clothes, collecting sufficient data is challenging. To mitigate data scarcity, we imported 3D clothes meshes into the simulation for data augmentation. Nevertheless, the limited availability of 3D mesh models restricted the model to extracting semantic keypoints from only four types of clothes. In this manuscript, we propose a novel semantic keypoint extraction pipeline for open-category clothes (Sec. 4). The proposed pipeline leverages large vision-language models and vision foundation models to overcome data scarcity and eliminate the need for training on labeled data. In this new pipeline, introducing a new clothes category requires only a prototype image—without any annotations—for autonomous semantic keypoint discovery (Sec. 4.1). This enhancement significantly improves the applicability and generalization of the proposed method.
A.2 Closed-loop Task Planning
In the conference version, task planning was performed in symbolic space—without conditioning on visual observations—and executed only once at the beginning of the manipulation task. This made the system sensitive to unexpected state changes, which are common when the initial clothes state is highly crumpled, as in flattening tasks. In this manuscript, we propose a closed-loop task planning pipeline (Sec. 5.2), where the task planning is conditioned on visual observations. After each sub-task is executed, the observation is updated, and the pipeline determines whether to replan. Additionally, each generated plan is verified before execution, and the system continuously checks whether the overall task has been completed. If the task is not completed, a new plan is generated based on the current observation. This closed-loop task planning pipeline enhances the system’s ability to handle unexpected state changes of clothes. Experimental results in Sec. 6.2 show that this enhancement significantly improves the system’s performance when the initial clothes state is crumpled.
A.3 Expanded Real-world Experimental Evaluation
Beyond the qualitative real-world results on four clothing categories presented in the conference version, we have extended our real-world experiments to include quantitative evaluations and analysis across a broader range of clothes categories, varying in size, shape, and material (Sec. 6.4). These new results further demonstrate the effectiveness of our proposed clothes manipulation method, which demands a higher level of generalization and robustness.
By introducing a novel open-category semantic keypoint extraction pipeline and a closed-loop task planning system, this manuscript presents substantial extensions over the conference paper.
Appendix B Prompts
Basic Skill Discovery Define basic skills for a dual-arm home-service robot, which is required to perform clothes manipulation tasks like folding, flattening, hanging, and placing.
Think about this step by step: 1. provides some examples of clothes manipulation tasks. 2. breaks down these tasks into basic actions. 3. summarizes basic actions and derive a set of basic skills.
Semantic Keypoint Discovery Select the ’semantic keypoints’ of clothes from ’candidate keypoints’ to solve the clothes manipulation tasks. Don’t ignore any semantic keypoints.
The input request is an image of the clothes, annotated with a set of visual marks: candidate keypoints: black dots marked as ’P[i]’ on the image, where [i] is an integer.
The definition of semantic keypoints: keypoints that are crucial for clothes manipulation tasks like folding, flattening, and hanging, these keypoints can be described with natural language. Please note: In the folding process, both picking and placing keypoints are ’semantic keypoints’. Don’t ignore any semantic keypoints. Don’t ignore the center if the center is the placing point. clothes often have symmetric structures, which can be useful guidance.
The response should be a list of ’semantic keypoints’, each semantic keypoint should be a dictionary that contains: keypoint: P1 language description: ‘left shoulder’ reason: ‘crucial for symmetry and folding the top part along with the left shoulder’
Coarse Region Matching Match the region in the test image that corresponds to the specified keypoint in the template image.
The input request contains: template image annotated with a keypoint. The keypoint is marked as a red dot. test image that is divided into several regions annotated with a set of visual marks.
The response should be the mark of the corresponding region in the following format:
The region corresponding to the red keypoint is ** **.
Sub-task Proposal Describe the dual-arm robot’s actions to solve the clothes manipulation tasks by: selecting contact points from annotated semantic keypoints. selecting basic skills from a pre-defined skill library.
You have the following basic skills to select: grasp(): pick up clothes parts using the grippers of one or both arms, the parameter can be one or two contact points. moveto(): transports the clothes or clothes parts to a target position, the parameter can be one or two contact points. release(): drops clothes parts by opening grippers. rotate(): rotates the clothes grasped by the gripper. pull(): stretches the clothes to flatten them.
The input request contains: instruction: a task described in natural language. image: a top-down photo of clothes, annotated with semantic keypoints labeled as P[i] (where i is the index). semantic keypoint dictionary: a dictionary mapping each P[i] to its language description.
You should determine whether the task is completed and the sequence of subtasks based on annotated keypoints. Each subtask includes the selected basic skill and contact points.
The response should be a dictionary in JSON form as follows:
{ “completed”: false, “subtasks”: [ “grasp(left sleeve)”, “moveto(center)”, “release()”, “grasp(right sleeve)”, “moveto(center)”, “release()”, “grasp(left shoulder, right shoulder)”, “moveto(left hem, right hem)”, “release()” ] }
I will provide some examples to demonstrate how to use these basic skills and the response format.
Appendix C Additional Clothes Details
This section presents the parameters of the clothes used in our experiment. The clothes cover various categories, shapes, textures, sizes, and materials. Each clothes is assigned a unique ID, and its size and material are also listed in the table. The size information indicates the width and height of the clothes when fully flat. In addition, we capture an RGB image of each clothes from a top-down view.
No. 1
Size 47 $\times$ 40 cm
Material
Cotton & Polyester
No. 2
Size 55 $\times$ 50 cm
Material
Polyester
No. 3
Size 60 $\times$ 43 cm
Material
Cotton & Polyester
No. 4
Size 40 $\times$ 30 cm
Material
Cotton
No. 5
Size 112 $\times$ 58 cm
Material
Wool
No. 6
Size 53 $\times$ 42 cm
Material
Polyester & Spandex
No. 7
Size 46 $\times$ 35 cm
Material
Cotton & Polyester
No. 8
Size 53 $\times$ 41 cm
Material
Cotton & Spandex
No. 9
Size 18 $\times$ 30 cm
Material
Cotton
No. 10
Size 53 $\times$ 41 cm
Material
Nylon
No. 11
Size 72 $\times$ 51 cm
Material
Polyester & Spandex
No. 12
Size 46 $\times$ 57 cm
Material
Cotton & Spandex
No. 13
Size 70 $\times$ 40 cm
Material
Polyester
No. 14
Size 101 $\times$ 42 cm
Material
Silk
No. 15
Size 98 $\times$ 35 cm
Material
Viscose